### v1

# Guided Transfer Learning for Discrete Diffusion Models

Abstract
Discrete diffusion models achieve strong performance across language and other discrete domains, providing a powerful alternative to autoregressive models. However, their strong performance relies on large training datasets, which are costly or risky to obtain, especially when adapting to new domains. Transfer learning is the natural way to adapt pretrained discrete diffusion models, but current methods require fine-tuning large diffusion models, which is computationally expensive and often impractical. Building on ratio-based transfer learning for continuous diffusion, we provide Guided Transfer Learning for discrete diffusion models (GTL). This enables sampling from a target distribution without modifying the pretrained denoiser. The same guidance formulation applies to both discrete-time diffusion and continuous-time score-based discrete diffusion, yielding a unified treatment. Guided discrete diffusion often requires many forward passes of the guidance network, which becomes impractical for large vocabularies and long sequences. To address this, we further present an efficient guided sampler that concentrates evaluations on planner-selected positions and top candidate tokens, thus lowering sampling time and computation. This makes guided language modeling practical at scale for large vocabularies and long sequences. We evaluate GTL on sequential data, including synthetic Markov chains and language modeling, and provide empirical analyses of its behavior. This is a pre-print.
1 Introduction
Diffusion models (ho2020denoisingdiffusionprobabilisticmodels; song2021scorebasedgenerativemodelingstochastic) have become the dominant approach for generative modeling in continuous domains like image, video, and audio (lovelace2023latentdiffusionlanguagegeneration; ho2022videodiffusionmodels; liu2023audioldmtexttoaudiogenerationlatent). More recently, discrete diffusion models for language have achieved competitive performance and, by denoising tokens in parallel and in any order, offer practical advantages for controllable infilling and editing compared to left-to-right autoregressive decoding (gong2025scalingdiffusionlanguagemodels; li2022diffusionlm). Among existing discrete diffusion approaches, masked diffusion models (MDMs) have demonstrated remarkable performance in tasks such as reasoning (nie2025scalingmaskeddiffusionmodels; zheng2025maskeddiffusionmodelssecretly), planning (ye2025autoregressiondiscretediffusioncomplex), and infilling (gong2025scalingdiffusionlanguagemodels), often performing on par with state-of-the-art auto-regressive language models (ALMs).
An open problem in the field is to develop transfer learning approaches that enable leveraging the incredible performance of large pre-trained diffusion models in domains with scarce data (ouyang2024transferlearningdiffusionmodels). In the continuous domain, recent works have explored transfer learning approaches based domain classifiers (ratio estimators) (zhong2025domainguidancesimpletransfer; ouyang2024transferlearningdiffusionmodels). Such methods bypass the need of updating millions of parameters, providing an alternative to adapter-based finetuning (moon2022finetuning; xie2023difffitunlockingtransferabilitylarge) and reinforcement learning (fan2023dpokreinforcementlearningfinetuning; clark2024directlyfinetuningdiffusionmodels) in scenarios of extreme scarcity.
In this paper, we tackle the problem of extending domain classifier-based sampling for discrete diffusion models. We first develop the theoretical analogues of (ouyang2024transferlearningdiffusionmodels), leading to a first naïve implementation of (8) (In supplementary materials, algorithm 2). However, we show that such naïve implementation requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per denoising step, scaling poorly with sequence length $L$ and vocabulary size $|\mathcal{V}|$, limiting its applicability to domains such as language modeling.
Motivated by these challenges, we introduce Guided Transfer Learning (GTL) for discrete diffusion models. Our method leverages a pretrained source discrete diffusion model and introduces a compact ratio network that estimates the density ratio between the target and source data distributions. We demonstrate that sampling from the target domain can be achieved solely by training this ratio-based guidance module, enabling effective target sampling without updating the denoiser. This substantially reduces the training cost compared to conventional fine-tuning approaches. We establish a closed-form solution for the sampling method, and we further extend this argument to continuous-time score-based models. This generalization allows our approach to cover the two dominant diffusion frameworks: score-based models (lou2024discretediffusionmodelingestimating; sun2023scorebasedcontinuoustimediscretediffusion) and discrete-time diffusion models (austin2023structureddenoisingdiffusionmodels), making the method broadly applicable to diffusion modeling. This GTL construction does not follow directly from continuous diffusion models (ouyang2024transferlearningdiffusionmodels), since score-matching diffusion is not defined for categorical discrete tokens and cannot be derived from the score-matching loss function (song2021scorebasedgenerativemodelingstochastic). We further introduce a guided sampling algorithm that reduces the number of guidance calls per denoising step by focusing computation on planner-selected tokens. This design improves efficiency. Previewing our results, GTL surpasses vanilla and fine-tuned diffusion across all data-scarcity regimes while training only  7% as many parameters (Fig. 1).
We summarize our contributions as follows:
  1. 1.
    We introduce GTL, an effective transfer learning framework for discrete diffusion models. GTL keeps the source denoiser fixed while training a lightweight ratio estimator on both source and target data to enable sampling from the target distribution. GTL achieves scalability by using an unmasking position for the top-$n_{\text{ratio}}$ candidates, dramatically reducing the per-step computational cost, without sacrificing performance in practice.
  2. 2.
    We provide a theoretical analysis and formal theorem of the proposed guided transfer learning method, demonstrating that ratio-based guidance correctly recovers the target reverse transition.
  3. 3.
    We validate GTL and the guided sampler on sequential data, including synthetic Markov chains and arXiv-abstract language modeling tasks.
Figure 1: GTL outperforms with  7% of the parameters. MAUVE ($\uparrow$) vs. fraction of target-domain training data (e.g., 100% = all 79,631 arXiv Physics abstracts; $1\%\approx 800$). Our Guided Transfer Learning (GTL, green) achieves the highest MAUVE across all data regimes while training only a 4.1M parameter ratio model outperforming both Vanilla (trained only on target data) and Fine-tuned (source-pretrained then fine-tuned on target) models. Points show mean, and error bars show standard error over 3×3 (train×sample) seeds. See Section 5.2 for the experimental setup and evaluation details.
2 Related Work
Discrete Diffusion Models
Discrete diffusion models (DDMs) are widely used for biological sequence generation (e.g., DNA, RNA, proteins) as well as language modeling (Sarkar2024; yu2025discretediffusionlargelanguage). Continuous‑noise variants that corrupt embeddings or latents are likely to suffer from embedding‑to‑token projection mismatch  (li2022diffusionlmimprovescontrollabletext; lovelace2023latentdiffusionlanguagegeneration). Beyond that, two widely adopted formulations operate entirely in discrete space, avoiding continuous mappings. First, simplified masked diffusion models such as MLDM and MD4 follow the D3PM framework, using discrete‑time transition matrices with a denoiser trained to invert the forward kernels (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion; austin2023structureddenoisingdiffusionmodels). Second, continuous‑time discrete diffusion estimates a discrete score, which represents the log‑ratio between adjacent‑timestep marginals. This induces a continuous-time Markov chain (CTMC) (lou2024discretediffusionmodelingestimating; meng2023concretescorematchinggeneralized; campbell2022continuoustimeframeworkdiscrete).
Guided Diffusion Models
In both continuous and discrete settings, guidance strategies like classifier‑based or classifier‑free are widely used. In the discrete case, classifier‑free guidance typically performs two forward passes at inference time, one conditional and one unconditional, while training a single large model on a mixture of conditional prompt–input pairs and unconditional data (huang2025ctrldiffboostinglargediffusion). Closer to our approach is classifier‑based guidance, which steers generation using a classifier $p(y\mid x_{t},t)$ while the denoiser is trained on the full corpus; sampling is biased toward the target label $y$ in the data distribution (schiff2025simpleguidancemechanismsdiscrete; nisonoff2025unlockingguidancediscretestatespace). In both discrete- and continuous-time formulations, this typically yields per-token factorized guidance, which incurs many computational steps, scaling with $\mathcal{O}(L|\mathcal{V}|)$ evaluations per denoising step. To mitigate this cost, (nisonoff2025unlockingguidancediscretestatespace) proposes a first-order Taylor approximation that reduces guidance evaluations to a forward and backward pass. TESS2 (tae2025tess2largescalegeneralist) instead applies gradient ascent on a reward with respect to the model logits during inference to improve response quality. Finally, the Energy‑based Diffusion Language Model (xu2025energybaseddiffusionlanguagemodels) defines an energy over predicted clean sequences and uses it to importance‑weight the denoising step, reducing the train–sample mismatch in discrete diffusion models.
Transfer Learning
Fine-tuning and transfer learning share the goal of adapting a model to a new task or domain. Fine-tuning is computationally expensive because it updates the pretrained denoiser’s weights on new data. Prior work on continuous diffusion models explores reinforcement learning (fan2023dpokreinforcementlearningfinetuning; clark2024directlyfinetuningdiffusionmodels) and adapter-based fine-tuning (moon2022finetuning; xie2023difffitunlockingtransferabilitylarge). In contrast, transfer learning for diffusion remains comparatively underexplored. Transfer Learning for Diffusion Models (TLDM) proposes a ratio-estimation technique for continuous score-based models, but demonstrates effectiveness on synthetic data and a one-dimensional heartbeat dataset, leaving applications to higher-dimensional and large-scale domains open.
3 Background
Diffusion models are generative models defined by a forward noising process and a learned reverse denoising process. In the reverse process, a denoiser $p_{\theta}$ is trained to remove noise from latent variables $\mathbf{z}_{t}$ to recover data $x$. The forward process generates $\mathbf{z}_{t}$ by injecting noise into clean data. We consider discrete latents: $\mathbf{z}_{t}\in\mathcal{V}$, the set of one-hot tokens over a vocabulary of size $N$, $\mathcal{V}\coloneqq\{\mathbf{z}\in\{0,1\}^{N}:\sum_{i=1}^{N}z_{i}=1\}\subset\Delta^{N-1},$ where $\Delta^{N-1}$ denotes the probability simplex over $N$ categories. Let $\text{Cat}(\cdot;\mathbf{\pi})$ be the categorical distribution over $N$ classes with the probability vector $\mathbf{\pi}\in\Delta^{N-1}$.
Discrete diffusion
Discrete Denoising Diffusion Probabilistic Models (D3PMs) (austin2023structureddenoisingdiffusionmodels) provide a flexible framework for diffusion over discrete state spaces. The forward Markov chain is $p(\mathbf{z}_{t}\mid\mathbf{z}_{s})=\text{Cat}\!\left(\mathbf{z}_{t};\,Q_{t\mid s}\mathbf{z}_{s}\right)$ where $Q_{t\mid s}\in\mathbb{R}^{N\times N}$ is a transition matrix whose $(i,j)$ entry is the probability of moving to state $i$ at time $t$ from state $j$ at time $s$. To fit the generative model $p_{\theta}$ to the data distribution over $\mathbf{x}_{0}$, we minimize the negative ELBO (NELBO). Different choices of $Q_{t\mid s}$ induce different corruption behaviors, e.g., uniform, absorbing, or discretized-Gaussian transitions.
Masked (absorbing-state) diffusion.
The absorbing-state variant or often called masked diffusion, replaces tokens with a special [MASK] token (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion). It can be viewed as an interpolation between the clean distribution and a noise prior over the vocabulary. For a clean token $\mathbf{x}_{0}$, the forward transition at time $t$ is
$p(\mathbf{z}_{t}\mid\mathbf{x}_{0})=\text{Cat}\!\left(\mathbf{z}_{t};\,\alpha_{t}\mathbf{x}_{0}+(1-\alpha_{t})\,\mathbf{m}\right),$ (1)
where $\alpha_{t}\!\in[0,1]$ controls the corruption schedule and $\mathbf{m}\in\mathcal{V}$ is one-hot vector at the special [MASK] token. Making use of the masking process, the backward posterior simplifies to:
$p(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})$
$=\begin{cases}\operatorname{Cat}\!\left(\mathbf{z}_{s};\,\mathbf{z}_{t}\right),&\text{if }\mathbf{z}_{t}\neq\mathbf{m},\\[6.0pt] \operatorname{Cat}\!\left(\mathbf{z}_{s};\,\dfrac{(1-\alpha_{s})\mathbf{m}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{0}}{1-\alpha_{t}}\right),&\text{if }\mathbf{z}_{t}=\mathbf{m}.\end{cases}$ (2)
Hence, when a token is unmasked ($\mathbf{x}_{t}\neq\mathbf{m}$) the step is deterministic: $\mathbf{x}_{s}=\mathbf{x}_{t}$. Otherwise, the posterior is a linear interpolation between the mask distribution and the clean token. This structure yields the simplified diffusion loss (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion):
$\mathcal{L}_{\text{diffusion}}$ $=\sum_{i=1}^{T}D_{\mathrm{KL}}\!\left(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\right)$ (3)
$=\sum_{i=1}^{T}\mathbb{E}_{q}\!\left[\frac{\alpha_{t(i)}-\alpha_{s(i)}}{1-\alpha_{t(i)}}\log\big\langle\mathbf{x}_{\theta}(\mathbf{z}_{t(i)}),\mathbf{x}_{0}\big\rangle\right],$ (4)
where $\mathbf{x}$ is the one-hot encoding of the clean data and $\langle\cdot,\cdot\rangle$ is the inner product. The network $\mathbf{x}_{\theta}$ predicts the clean token distribution from the noisy input $\mathbf{z}_{t(i)}$.
Guidance
Diffusion models provide strong controllability through both classifier-based and classifier-free guidance. These two approaches correspond to different ways of expressing the backward generative process conditioned on $y$. In this work, we focus on the classifier-based variant, which allows us to rearrange the conditional generative process using Bayes’ rule as $p^{\gamma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},y)=p(y\mid\mathbf{z}_{s},\mathbf{z}_{t})^{\gamma}\,p(\mathbf{z}_{s}\mid\mathbf{z}_{t})/Z$, with $Z$ being a normalizing constant.
For the extension to sequences of tokens $\mathbf{z}_{s}^{(1:L)}$ of length $L$, we assume that the distribution $p^{\gamma}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)},y\right)$ factorizes independently across tokens. This leads to the following formulation:
$p^{\gamma}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)},y\right)$
$=\prod_{\ell=1}^{L}\frac{p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}p\!\left(y\mid\tilde{\mathbf{z}}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}$ (5)
Here, $\mathcal{V}$ denotes the vocabulary set, and $p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)$ represents the classifier probability given the sequence $\mathbf{z}_{t}^{(1:L)}$ with the token $\mathbf{z}_{s}$ placed at position $\ell$. However, this formulation comes with the drawback of requiring $\mathcal{O}(L\cdot|\mathcal{V}|)$ forward passes through the classifier model.
Problem Formulation
Let $\mathcal{X}$ denote the data space, with source domain distribution $p_{X}$ and target domain distribution $q_{X}$. We are given source samples $\mathcal{S}=\{\mathbf{x}_{i}^{(1:L)}\}_{i=1}^{m}\sim p_{X}$ and target samples $\mathcal{T}=\{\mathbf{x}^{(1:L)}_{i}\}_{i=1}^{n}\sim q_{X}$. The goal of transfer learning is to train a network $p_{\theta}$ on $\mathcal{S}$ and adapt it to $\mathcal{T}$ to improve performance in the target domain. In practice, $n\ll m$, creating an imbalance between source and target data. A common strategy is to freeze most pretrained layers, extract features, and train new task-specific layers on $\mathcal{T}$. Fine-tuning relaxes this by unfreezing parameters, allowing the model to adapt more flexibly to the target distribution. In our setting, we are given a pretrained generative model $p_{\theta}$ trained on the source distribution $p_{X}$, and we aim to transfer its generative process to the target distribution $q_{X}$, yielding a target-domain model $q_{\psi}$. Note that the background section is introduced for the source domain diffusion.
4 Guided Transfer Learning for DDMs
We propose Guided Transfer Learning for Discrete Diffusion Models, a framework that enables transfer learning for discrete diffusion models. We first formalize the problem setup, then present the mathematical formulation of our method, and finally introduce a new sampling algorithm that leverages our results to make transfer learning broadly applicable across datasets.
Given the problem statement and access to both domains, one could fine-tune the source network $p_{\theta}$ by maximizing (4), thereby optimizing the entire parameter set. For large language models, this is often costly and slow due to the scale of parameters. Moreover, when the target domain provides far fewer samples, fine-tuning alone rarely yields a strong generative model in terms of diversity and quality.
Our method is based on the following theorem, which proposes a different approach to this problem: it leverages a pretrained source-domain diffusion model together with a ratio estimator between the two domains. We assume the two diffusion models share the same forward process $p(\mathbf{z}_{t}\mid\mathbf{x}_{0})=q(\mathbf{z}_{t}\mid\mathbf{x}_{0})$ for all $t$, that the density ratio $q(\mathbf{x}_{0})/p(\mathbf{x}_{0})$ is well defined.
Relation to TLDM. TLDM (ouyang2024transferlearningdiffusionmodels) proves a transfer result for continuous, score-based diffusion and is not directly applicable to discrete state spaces. The theorem is stated in $\mathbb{R}^{d}$ and relies on score matching with differentiable densities, requiring gradients such as $\nabla_{x_{t}}\log q_{t}(x_{t})$. In the discrete setting (categorical tokens), such Euclidean scores are undefined. Beyond this, the modeling and training objectives differ: continuous diffusion injects Gaussian noise and optimizes score-matching Mean Squared Error, whereas discrete diffusion uses categorical transition matrices (e.g., absorbing/uniform) and is trained by minimizing KL between categorical true posteriors $q(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})$ and $q_{\psi}(\mathbf{z}_{s}\mid\mathbf{z}_{t})$.
Theorem 1
Let $p$ (source) and $q$ (target) be two diffusion models that share the same forward process, $p(\mathbf{z}_{t}\mid\mathbf{x}_{0})=q(\mathbf{z}_{t}\mid\mathbf{x}_{0})$ for all $t$. The reverse target domain network $q_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ is defined by:
$\underset{\psi}{\arg\min}\;\mathbb{E}_{q}\!\left[\sum_{i=1}^{T}D_{\mathrm{KL}}\!\Bigl(q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\Big\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\Bigr)\right]$ (6)
Then we have
$q_{\psi^{\star}}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})=\frac{p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}{\displaystyle\sum_{\mathbf{\tilde{z}}_{s(i)}}p(\mathbf{\tilde{z}}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{\tilde{z}}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}$ (7)
In (7), $p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\approx p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ is the trained denoiser on the source domain, and the expectation factor $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\tfrac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right]$ is the guidance ratio. Thus, the target diffusion process can be approximated with two networks $\psi^{\star}=[\theta,\phi]$. The proof of this theorem can be found in the supplementary materials. The advantage of this approach is that it does not require fine-tuning the pretrained discrete diffusion model. Instead, the main computational cost is shifted to training a domain classifier using both domains as data. Importantly, the domain classifier can be much smaller than the discrete diffusion model. Moreover, density ratio estimation provides more flexibility in regularization when dealing with highly imbalanced data. It only requires distinguishing domains rather than modeling the full target distribution, which makes the training dynamics between the discriminative and generative models more stable.
The result in (ouyang2024transferlearningdiffusionmodels) is derived only for score-based diffusion in continuous data domains. But our formulation in (7) offers two advantages: (i) it permits flexible definitions of the forward and reverse discrete diffusion processes, and (ii) the guidance term can be trained under arbitrary noise processes, in either continuous- or discrete-time settings.
In the supplementary materials (Theorem 3), we extend the result to the score-based, continuous-time diffusion setting over discrete state spaces. The key mathematical step is to trace back through to the ELBO formulation used in (6). For the masked (absorbing-state) diffusion training objective, this link is directly shown in Equation (4).
Extension to Sequence
Leveraging our theoretical results, (7) closely parallels (5). Instead of a classifier $p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)$, we use a ratio network $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\tfrac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right]$. To extend this to sequences, define $\mathbf{z}_{t,s^{\ell}}^{(1:L)}\coloneqq[\,\mathbf{z}_{t}^{(1:\ell-1)},\,\mathbf{z}_{s}^{(\ell)},\,\mathbf{z}_{t}^{(\ell+1:L)}\,]$, i.e., the sequence identical to $\mathbf{z}_{t}^{(1:L)}$ except at position $\ell$, where the token is replaced by $\mathbf{z}_{s}^{(\ell)}$. This yields the following factorized, ratio-guided conditional (with guidance strength $\gamma\geq 0$):
$q_{\theta,\phi}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)}\right)$
$=\prod_{\ell=1}^{L}\frac{\Big[r_{\phi}\!\big(\mathbf{z}_{t,s^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\displaystyle\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}\Big[r_{\phi}\!\big(\mathbf{z}_{t,\tilde{s}^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}.$ (8)
A naive integration of this guidance into the sampling process incurs a computational complexity of $\mathcal{O}(|\mathcal{V}|\cdot L)$ per denoising step, since we must evaluate $r_{\phi}$ for all sequences obtained by replacing every possible token in $\mathcal{V}$ at each position $L$ in the denominator. Prior work (schiff2025simpleguidancemechanismsdiscrete; nisonoff2025unlockingguidancediscretestatespace) targets settings with modest vocabulary sizes and sequence lengths, where this cost is manageable. In our experiments with the Markov chain (Section 5.1) it is also not prohibitive. But for text applications, the cost becomes significant. (nisonoff2025unlockingguidancediscretestatespace) proposes a first-order Taylor approximation that reduces evaluations, but it is accurate primarily in the high-noise regime and degrades as $t$ decreases.
4.1 Scalability via Planner Sampling
In this section, we tackle the main bottleneck: integrating guidance efficiently at inference time. The gains from a set of simple techniques are summarized in Table 1.
Method Per-Step Complexity Number of Steps ($T$)
Ancestral Sampling $\mathcal{O}(|\mathcal{V}|\cdot L)$ $T$
+ Caching $\mathcal{O}(|\mathcal{V}|\cdot L)$ $T^{\prime}\leq\min(L,T)$
+ Top-$n_{\text{ratio}}$ $\mathcal{O}(n_{\text{ratio}}\cdot L)$ $T^{\prime}\leq\min(L,T)$
Planner Sampling
+ Top-$n_{\text{ratio}}$ $\mathcal{O}(n_{\text{ratio}})$ $L$
Table 1: Computational cost per denoising step and total number of steps.
Ancestral Sampling
In the naive implementation, at all denoising steps, we compute logits from the denoiser and from the guidance network with $\mathcal{O}(|\mathcal{V}|\cdot L)$ complexity. Following (sahoo2024simpleeffectivemaskeddiffusion), in masked diffusion, we can remove explicit time conditioning because the number of masked tokens in the sequence already implies the timestep. This is not possible for uniform noise, which mixes token identities. We therefore cache logits across steps. If a step does not change the sequence, we reuse the logits from the previous step. The number of sampling steps where the network is called then becomes upper-bounded by the sequence length, since in the worst case, every position unmasks in a different step.
In addition, we introduce a top-$n$ style pruning mechanism for the ratio network. Specifically, we assume that (i) the denoiser distribution $p_{\theta}(\cdot\mid z_{t})$ is sufficiently concentrated, so that the true target distribution $q$ differs from $p$ only within the top-$n_{\text{ratio}}$ candidates, and (ii) the density ratio $r(z)=q(z)/p(z)$ is bounded, in the sense that there exists a constant $C<\infty$ such that for all tokens within the top-$n_{\text{ratio}}$ set of $p_{\theta}$, we have $\frac{q(z)}{p(z)}\leq C$. Outside this set, where $p(z)$ is already very small, no such bound is required. Consequently, we evaluate the ratio term only for the top $n_{\text{ratio}}$ candidates proposed by the denoiser at each position. This results in position-specific candidate sets in the denominator of equation (8) and reduces the computational cost of ratio evaluations to $\mathcal{O}(n_{\text{ratio}}\cdot L)$ with $n_{\text{ratio}}\leq|\mathcal{V}|$. In the experiments, the combination of top-$n_{\text{ratio}}$ selection and the guidance weight $\gamma$ can misallocate probability mass, which may cause the diffusion process to collapse into a fully masked sequence at the final step. Thus, a key assumption is that the probability of remaining a masked token must be equal between the target and source distributions. As a result, we enforce a position-independent normalization which implies the renormalized update over non-mask tokens, thus stabilizing the denoising diffusion process. Further details of this stabilizer are presented in the supplementary materials.
Algorithm 1 Ratio-Guided Denoise Step $(\mathbf{z}_{t},t,\Delta t)$.
0: noisy sequence $\mathbf{z}_{t}\in\mathcal{V}^{L}$, denoiser $p_{\theta}$, ratio net $r_{\phi}$, guidance schedule $\gamma(\cdot)$, planner $\rho_{\vartheta}$
[…]
4:for all $v\in\mathcal{N}$ do
5:  $\tilde{\mathbf{z}}\leftarrow(\mathbf{z}_{t}[\ell]\leftarrow v)$
6:  $\log r[\ell,v]\leftarrow r_{\phi}(\tilde{\mathbf{z}},\sigma_{t})$ {Ratio scores}
7:end for
[…]
13:return $\mathbf{z}_{s}$
Planner Sampling
We also apply our method to text in Section 5.2, where $n_{\text{ratio}}\ll\lvert\mathcal{V}\rvert$, yet sampling remains slow and costly mainly because of the $L$ in the complexity. To address this, we propose a planner-based algorithm. The planner $\rho_{\vartheta}$ is trained to predict the next denoising position in the masked sequence. Concretely, we replace the Gillespie process with $\tau$-leaping by a planner schedule that unmasks exactly one position per step, thereby fixing the number of denoising steps to $L$. In many implementations, $T\approx L$, so enforcing a hard budget of $L$ steps does not materially change the step count, while reducing guidance evaluations and wall-clock time. Pseudocode is provided in Algorithm 1. An important achievement here is that the loop in lines 4 to 7 can now be parallelized by vectorization on a single GPU. As a result, only one forward pass of the three networks is required.
For training the planner, we adopt a simplified training algorithm following (peng2025pathplanningmaskeddiffusion; liu2025thinkgeneratediscretediffusion). For ratio training, we use a discrete data training loss function building on the work of (ouyang2024transferlearningdiffusionmodels). Further details of the loss function and training algorithm are provided in the supplementary materials.
5 Experiments
In this section, we present empirical results on the effectiveness of our method, Guided Transfer Learning for discrete diffusion models (GTL). We provide evidence across two applications showing that our method can successfully sample from the target distribution, in comparison to a vanilla diffusion model and standard finetuning. The first section presents a proof-of-concept on a synthetic dataset based on learning Markov chain data. The second subsection reports results on language modeling using arXiv abstracts.
5.1 Synthetic Dataset
Refer to caption (a) SourceRefer to caption (b) Target
n Vanilla Discrete Diffusion Finetuned Discrete Diffusion GTL
1000 Refer to caption Refer to caption Refer to caption
100 Refer to caption Refer to caption Refer to caption
20 Refer to caption Refer to caption Refer to caption
Figure 2: True source and target (left). Estimated target transition matrices for varying $n$ across three methods (right).
Experimental Setup
We construct two Markov chains with distinct transition matrices. Let the probability state vector be $\mathbf{\pi}_{i}\in\Delta^{N-1}$ over a vocabulary of size $N=5$. The transition matrix of the target domain is $\mathbf{T}\in\mathbb{R}^{N\times N}$ and that of the source domain is $\mathbf{S}\in\mathbb{R}^{N\times N}$. We sample sequences of length $20$ from the target domain using the update rule $\mathbf{\pi}_{i}=\mathbf{T}\,\mathbf{\pi}_{i-1}$ with the initial state $\mathbf{\pi}_{0}$ set to the uniform vector. Sampling from the source domain proceeds in the same way.
Implementation
We use the sampling process of (sahoo2024simpleeffectivemaskeddiffusion) for the source, finetuned, and target discrete diffusion models. The backbone is a Diffusion Transformer (DiT) (peebles2023scalablediffusionmodelstransformers) with about five million learnable parameters. The ratio estimator and the classifier use about half as many parameters with a final mean and fully connected layer that maps the output to a scalar. The exact implementation appears in our codebase. For guided sampling, we follow the ancestral sampling method. Since the vocabulary is small we set $n_{\text{ratio}}=N$ and use $20$ sampling steps. We train a masked diffusion model on the source domain for $30$ epochs. We then compare a vanilla masked diffusion model trained directly on the target samples (60 epochs), a finetuned diffusion model (90 epochs), and our GTL method (60 epochs). The finetuned model and our method share the same source diffusion model.
Results
We consider $m=10000$ source samples and $n\in\{1000,100,20\}$ target samples. We estimate the transition matrix by sampling $4096$ sequences, counting transitions between states, and normalizing by the total count. Figure 2 presents an example where $\mathbf{S}$ has diagonal entries $0.1$ and $\mathbf{T}$ has diagonal entries $0.8$. We observe that the vanilla target diffusion model and the finetuned model both degrade in performance when fewer target samples are available.
$n_{\text{tgt}}$ Vanilla (Target-only) Finetuned GTL
[…]
20 $0.5842\pm 0.0244$ $0.4004\pm 0.0983$ $\mathbf{0.3621}\pm 0.0478$
Table 2: Comparison of baseline methods and best guided results (mean $\pm$ standard error; n=3n{=}3 seeds) across target sample sizes for $(\text{diag}_{\text{src}}=0.1,\text{diag}_{\text{tgt}}=0.8)$. Reported values are the average Kullback–Leibler divergence between true and estimated transition matrices; lower is better.
Table 2 provides a direct comparison, obtained by running this example three times with different random seeds. We observe the following behaviors: (i) the diffusion model trained directly on the target dataset loses performance as the number of training samples decreases, and (ii) our method, compared to traditional finetuning, achieves better results even with fewer trainable parameters, while showing a similar trend in performance across different numbers of training samples.
5.2 Language Modeling
Refer to caption Refer to caption Refer to caption
(a) Effect of $n_{\text{ratio}}$ at $\gamma=5$. (b) Effect of low $\gamma$ at fixed $n_{\text{ratio}}$ (c) MAUVE vs. high $\gamma$
[…]
In this section, we apply our method to text data. We use the arXiv abstracts dataset (arxiv_org_submitters_2024), which contains 1.7 million arXiv articles with titles, authors, categories, and abstracts. We focus on Computer Science, Mathematics, and Physics. For data preparation, we concatenate all text and split it into segments of 512 tokens using the bert-base-uncased tokenizer with vocabulary size N=30,522N=30{,}522. This results in 285,946285{,}946 samples in Computer Science, 176,831176{,}831 in Mathematics, and 79,63179{,}631 in Physics. For each category we randomly select 10% of the samples for validation. We ensure that the three category subsets are disjoint. We define the union of Computer Science and Mathematics as the source domain and Physics as the target domain.
We run two types of experiments. First, we study the behavior of the two key parameters $\gamma$ and $n_{\text{ratio}}$. Second, we compare three settings on the target domain. The first is a vanilla discrete diffusion model trained on the target data. The second is a finetuned version of the source model. The third is our guided transfer method.
Implementation and Evaluations
We train the source diffusion model for 100,000 gradient steps using the DiT architecture (peebles2023scalablediffusionmodelstransformers) with 59.8M trainable parameters. A log–linear noise scheduler is used for both training and sampling. The finetuned model on the target domain is trained for 10,000 gradient steps. For our guided method, we initialize a ratio model with 4.1M parameters and train it for 5,000 gradient steps. Additional training details appear in the supplementary materials.
For each method we generate 128 sequences of length 512. We evaluate quality with MAUVE (pillutla2021mauvemeasuringgapneural), which captures the tradeoff between perplexity and diversity. Although generative perplexity (Gen. PPL) is often reported, we observed, consistent with prior work (wang2025remaskingdiscretediffusionmodels), that Gen. PPL can remain low even when the text has reduced utility, such as repeated words. In those cases, MAUVE is more sensitive and provides a clearer measure of quality. We also evaluate the domain of each sampled sequence using an independently trained classifier. Scores near zero indicate the target domain and scores near one indicate the source domain. The classifier follows the setup used for the ratio model in the supplementary materials with label smoothing set to $0.05$.
Behavior of $\gamma$ and $n_{\text{ratio}}$
We study the sensitivity of GTL to the guidance weight $\gamma$ and the candidate budget $n_{\text{ratio}}$. For each configuration, we train three independent source models and three ratio models using different random seeds. The source domain is defined as Computer Science $\cup$ Mathematics plus an $r$-fraction of Physics abstracts, and the target domain is the remaining (1r)(1-r) fraction of Physics. Formally, for $r\in[0,1]$, let $\mathcal{P}_{r}$ be a random $r$-subset of Physics (sampled without replacement). We set $\mathcal{S}=\text{CS}\cup\text{Math}\cup\mathcal{P}_{r}$ and $\mathcal{T}=\mathcal{P}\setminus\mathcal{P}_{r}$. We set $r\in\{0,0.05,0.25\}$, thus in the case of $r=0$ we have two distinct datasets. This setup simulates the common scenario where pretraining large corpora already includes a small portion of target-domain text. In Figure 3, the behavior of these parameters is illustrated. Intuitively, smaller candidate budgets $n_{\text{ratio}}$ introduce higher sampling error, whereas increasing $n_{\text{ratio}}$ monotonically reduces the domain-classifier score, yielding more target-like samples, as shown in panel (a). In Figure 3(b), we observe that the guidance score is nontrivial: a moderate $\gamma\approx 3\!-\!5$ consistently pushes the representations toward the target across all $r$. Panel (c) highlights two key effects: (i) excessive guidance can degrade performance when there is no overlap between source and target data, and (ii) when a small portion of target text is included in the source domain ($r=0.05,0.25$), MAUVE remains high for larger $\gamma$ values and decays more smoothly. This indicates that a higher $r$ improves robustness. Overall, the results suggest that $n_{\text{ratio}}\approx 256\!-\!512$ and $\gamma\approx 4\!-\!6$ achieve a favorable balance, providing strong target alignment with minimal quality loss.
Method Comparison and Sensitivity to Data Fraction
We compare three methods: a fine-tuned discrete diffusion model trained on the target domain after pretraining on the source domain, a vanilla discrete diffusion model trained directly on the target domain, and our Guided Transfer Learning (GTL) method with the ratio estimator trained across target-data fractions of 100%, 30%, 5%, and 1%. Each method is trained with three random seeds, and sampling is performed with three seeds as well (nine runs per cell). Furthermore, each seed corresponds to a distinct subset of the target data. Figure 1 reports the behavior of the models across different fractions of the target training data. The observed trends align with those in Table 2. In the smallest case, we only use 796 samples of Physics abstracts. For finetuned and vanilla discrete diffusion, the number of epochs over the training set increases up to 400 when the fraction is smallest. The ratio model decreases its MAUVE score with decreasing fractions, but less dramatically than the other methods. Thus, our method demonstrates greater stability under limited target-domain data. In contrast, the finetuned and vanilla discrete diffusion models show a stepwise drop in performance as the fraction decreases. The GTL remains relatively stable, except in the 1% case, where we also observe a sharp decline.
6 Conclusion
Adapting discrete diffusion models to target domains with limited data is often costly. To address this, we propose a guided transfer learning approach for discrete diffusion models that introduces a small, learned ratio network, allowing the training of significantly fewer parameters. Furthermore, we present a guidance-based sampling algorithm that enables generation with larger vocabularies and longer sequences, making guided language modeling with discrete diffusion models feasible. We demonstrate the effectiveness of our method on synthetic Markov chains and arXiv abstracts (language modeling), particularly when the target dataset is small relative to the source.
Limitations
Overall, the proposed method is promising and applicable to a wide range of tasks involving domain shifts. However, it requires access to both source and target domains. Future work could explore conditional transfer for prompt-based or attribute-controlled generation. It would also be interesting to evaluate the learned ratio on quality-filtered tasks, such as reducing toxicity in language modeling.
7 Supplementary Materials for Guided Transfer Learning for Discrete Diffusion Models
Appendix A Proof of Theorem 1
For the proof, we use the formulation of importance reweighting as described in the following lemma.
Lemma 2 (Importance-weighted reformulation)
Fix a timestep $i$ and define $r(\mathbf{x}_{0})=q(\mathbf{x}_{0})/p(\mathbf{x}_{0})$. If the source and target share the same forward process, then the loss function can be rewritten as:
$\mathcal{L}_{i}(\psi)$ $=\mathbb{E}_{q}\bigl[D_{\mathrm{KL}}\!\bigl(q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)\bigr]$ (9)
$=\mathbb{E}_{\mathbf{x}_{0}\sim p\;z_{t(i)}\sim p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}\Bigl[r(\mathbf{x}_{0})D_{\mathrm{KL}}\!\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)\,\Bigr].$ (10)
Proof of Lemma 2
$\mathbb{E}_{q}\!\left[\mathcal{L}_{\text{diffusion}}\right]$ $=\sum_{i=1}^{T}\mathbb{E}_{q}\!\Bigl[D_{\mathrm{KL}}\!\bigl(q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)\Bigr]$ (11)
$=\sum_{i=1}^{T}\sum_{\mathbf{x}_{0}}\sum_{\mathbf{z}_{t(i)}}q(\mathbf{x}_{0},\mathbf{z}_{t(i)})D_{\mathrm{KL}}\!\bigl(q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)$ (12)
$=\sum_{i=1}^{T}\sum_{\mathbf{x}_{0}}\sum_{\mathbf{z}_{t(i)}}q(\mathbf{x}_{0})\,q(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})D_{\mathrm{KL}}\!\bigl(q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)$ (13)
$=\sum_{i=1}^{T}\sum_{\mathbf{x}_{0}}\sum_{\mathbf{z}_{t(i)}}{\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}}p(\mathbf{x}_{0})\,p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})D_{\mathrm{KL}}\!\bigl(q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)$ (14)
$=\sum_{i=1}^{T}\mathbb{E}_{\mathbf{x}_{0}\sim p}\;\mathbb{E}_{\mathbf{z}_{t(i)}\sim p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}\Bigl[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}D_{\mathrm{KL}}\!\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)\,\Bigr]$ (15)
To conclude Lemma 2, we leverage the fact that the source and target domain share the same forward kernel in Equation 14 to 15 $q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=\frac{q(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\cdot q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})}{q(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$.
Now, we use Lemma 2:
$\mathcal{L}_{i}(\psi)$ $=\mathbb{E}_{\mathbf{x}_{0}\sim p}\;\mathbb{E}_{\mathbf{z}_{t(i)}\sim p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}\Bigl[\underbrace{\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}}_{r(\mathbf{x}_{0})}D_{\mathrm{KL}}\!\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)\,\Bigr]$ (16)
$=\sum_{\mathbf{x}_{0}}\sum_{\mathbf{z}_{t(i)}}\underbrace{p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})\;p(\mathbf{x}_{0})}_{p(\mathbf{x}_{0}\mid\mathbf{z}_{t(i)})\;p(\mathbf{z}_{t(i)})}r(\mathbf{x}_{0})D_{\mathrm{KL}}\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\;\|\;q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)$ (17)
$=\sum_{\mathbf{z}_{t(i)}}p(\mathbf{z}_{t(i)})\sum_{\mathbf{x}_{0}}\underbrace{p(\mathbf{x}_{0}\mid\mathbf{z}_{t(i)})r(\mathbf{x}_{0})}_{\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})}D_{\mathrm{KL}}\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\;\|\;q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)$ (18)
$=\sum_{\mathbf{z}_{t(i)}}p(\mathbf{z}_{t(i)})\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\;D_{\mathrm{KL}}\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\;\|\;q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)$ (19)
Now, we take the derivative in regard to $\psi$:
$\qquad\qquad 0=\frac{\partial}{\partial\psi}\mathcal{L}_{i}(\psi)$ (20)
$\Longrightarrow$ $\qquad\qquad 0=\frac{\partial}{\partial\psi}\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\;D_{\mathrm{KL}}\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\;\|\;q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)$ (21)
To solve this, we set up a Lagrangian equation by using the constraint $\sum_{\mathbf{z}_{s(i)}}q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})=1$
$\mathcal{L}(\psi,\lambda)\;=\;$ $\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\sum_{\mathbf{z}_{s(i)}}p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\bigl[\log p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})-\log q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr]$
$+\;\lambda\!\Bigl(\sum_{\mathbf{z}_{s(i)}}q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})-1\Bigr)$ (22)
Stationary condition for each $\mathbf{z}_{s(i)}$:
$\frac{\partial\mathcal{L}}{\partial\psi}=-\frac{\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})}{q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})}\;+\;\lambda=0\quad$ (23)
$\Longrightarrow\quad{q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})=\frac{\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})}{\lambda}}$ (24)
Now we determine $\lambda$ from the constraint
$1$ $=\sum_{\mathbf{z}_{s(i)}}q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})=\sum_{\mathbf{z}_{s(i)}}\frac{1}{\lambda}\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$ (25)
$\;\Longleftrightarrow\qquad\lambda$ $=\sum_{\mathbf{z}_{s(i)}}\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$ (26)
Putting all together:
$q_{\psi^{\star}}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ $=\frac{\displaystyle\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})}{\displaystyle\sum_{\mathbf{\tilde{z}}_{s(i)}}\displaystyle\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})p(\mathbf{\tilde{z}}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})}$ (27)
$=\frac{p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}}{\displaystyle\sum_{\mathbf{\tilde{z}}_{s(i)}}p(\mathbf{\tilde{z}}_{s(i)}\mid\mathbf{z}_{t(i)})\frac{q(\mathbf{\tilde{z}}_{s(i)})}{p(\mathbf{\tilde{z}}_{s(i)})}}$ (28)
$={\frac{p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}{\displaystyle\sum_{\mathbf{\tilde{z}}_{s(i)}}p(\mathbf{\tilde{z}}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{\tilde{z}}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}}$ (29)
For 27 to 28 we show that $\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}$:
$\displaystyle\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$ $=\displaystyle\sum_{\mathbf{x}_{0}}p(\mathbf{x}_{0}\mid\mathbf{z}_{t(i)})\,\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$ (30)
$=\displaystyle\sum_{\mathbf{x}_{0}}p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})\,\frac{p(\mathbf{x}_{0})}{p(\mathbf{z}_{t(i)})}\,\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\,\frac{p(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\cdot p(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})}{p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}$ (31)
$=\frac{p(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})}{p(\mathbf{z}_{t(i)})}\displaystyle\sum_{\mathbf{x}_{0}}q(\mathbf{x}_{0})p(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})=p(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{t(i)})}$ (32)
For 28 to 29 we show that $\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}=\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\bigl[q(\mathbf{x}_{0})/p(\mathbf{x}_{0})\bigr]$:
$\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}$ $=\frac{\sum_{\mathbf{x}_{0}}q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})\,q(\mathbf{x}_{0})}{p(\mathbf{z}_{s(i)})}$ (33)
$=\frac{\sum_{\mathbf{x}_{0}}q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})\,p(\mathbf{x}_{0})\;r({\mathbf{x}_{0}})}{p(\mathbf{z}_{s(i)})}=\sum_{\mathbf{x}_{0}}\frac{p(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})\,p(\mathbf{x}_{0})}{p(\mathbf{z}_{s(i)})}\;r({\mathbf{x}_{0}})$ (34)
$=\sum_{\mathbf{x}_{0}}p(\mathbf{x}_{0}\mid\mathbf{z}_{s(i)})\;r({\mathbf{x}_{0}})=\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\bigl[q(\mathbf{x}_{0})/p(\mathbf{x}_{0})\bigr]$ (35)
Appendix B Guided $\tau$-Sampling
Algorithm 2 Naive Guided Denoise Step $(\mathbf{z}_{t},\ t,\ \Delta t)$
0: noisy sequence $\mathbf{z}_{t}\in\mathcal{V}^{L}$, time $t$, step $\Delta t>0$, denoiser $p_{\theta}$, ratio net $r_{\phi}$, mask index $m$, guidance schedule $\gamma(\cdot)$, noise schedule $\sigma(\cdot)$
0: updated sequence $\mathbf{z}_{s}$
[…]
5:for all $\ell\in E$ do
6:  for all $v\in\mathcal{V}$ do
7:   $\tilde{\mathbf{z}}\leftarrow\mathbf{z}_{t}$
8:   $\tilde{\mathbf{z}}^{(\ell)}\leftarrow v$
9:   $\log r[\ell,v]\leftarrow\log r_{\phi}(\tilde{\mathbf{z}},\ \sigma_{t})$
10:  end for
11:  $\log q^{\text{guided}}[\ell,:]\leftarrow\log x_{\theta}[\ell,:]+\gamma(t)\,\log r[\ell,:]$
12:  $\pi^{(\ell)}\leftarrow\operatorname{softmax}\!\big(\log q^{\text{guided}}[\ell,:]\big)$
13:  $z_{s}^{(\ell)}\sim\operatorname{Cat}\!\big(\pi^{(\ell)}\big)$
14:end for
15:
16:return $\mathbf{z}_{s}$
In the experiments, we observe that the combination of top-$n_{\text{ratio}}$ selection and the guidance weight $\gamma$ can misallocate probability mass, potentially causing the diffusion process to collapse into a fully masked sequence at the final step. For instance, when the top-$n_{\text{ratio}}$ is small, the denoiser tends to propose tokens that are highly likely to belong to the source domain, which then receive negative logits from the guidance ratio network. This effect is especially pronounced in high-noise regions, where all tokens experience a reduction in their probability of being unmasked, though some are affected more than others. Consequently, the overall probability of unmasking decreases, while the probability of remaining masked is incorrectly increased. This leads to the issue of ending up with a higher number of masked positions in the final step. To mitigate this problem, we apply a position-independent normalization:
$q_{\theta,\phi}^{\gamma}(\mathbf{z}_{s}=\mathbf{m}\mid\mathbf{z}_{t})=p(\mathbf{z}_{s}=\mathbf{m}\mid\mathbf{z}_{t})=p_{\mathbf{m}},\quad\quad\sum_{{}_{{\mathbf{v}}\in\mathcal{V}\setminus\{\mathbf{m}\}}}q_{\theta,\phi}^{\gamma}(\mathbf{z}_{s}=\mathbf{v}\mid\mathbf{z}_{t})=1-p_{\mathbf{m}},$
which implies the renormalized update over non-mask tokens, thus stabilizing the denoising diffusion process.
$q_{\theta,\phi}^{\gamma}(\mathbf{z}_{s}=\mathbf{v}\mid\mathbf{z}_{t})=\frac{(1-p_{\mathbf{m}})\,p_{\theta}(\mathbf{z}_{s}=\mathbf{v}\mid\mathbf{z}_{t})\,r_{\phi}(\mathbf{z}_{s}=\mathbf{v})^{\gamma}}{\displaystyle\sum_{\tilde{\mathbf{v}}\in\mathcal{V}\setminus\{\mathbf{m}\}}p_{\theta}(\mathbf{z}_{s}=\tilde{\mathbf{v}}\mid\mathbf{z}_{t})\,r_{\phi}(\mathbf{z}_{s}=\tilde{\mathbf{v}})^{\gamma}}\quad\text{for all }\mathbf{v}\in\mathcal{V}\setminus\{\mathbf{m}\}.$
We use this algorithm in the first experiments section.
Appendix C Training Details
The Code will be released upon acceptance. In this section, we briefly describe the training setup for reproducibility. All training networks share the same learning schedule, which includes cosine decay with warmup and a high learning rate of $3\times 10^{-4}$. The Adam optimizer is used, and we apply a dropout rate of 0.1. The batch size is set to 256 across all training algorithms. The time-independent and time-dependent Classifier is trained with 0.1 label smoothing and 4000 gradient steps. Importantly, the classifier is trained with binary labels, where the target domain is labeled as zero and the source domain as one. The ratio network is trained with $\lambda$ set to 0.1, as outlined in the following pseudo-algorithm 3.
Algorithm 3 Training loop for the ratio network $r_{\phi}(x_{t},t)$
[…]
1:repeat
2:  Sample mini‑batch $x_{0}$ of size $b$ from $\mathbf{p}$ and draw $t\sim\text{Uniform}(0,1)$
3:  $x_{t}\leftarrow\textsc{Corrupt}(x_{0},\sigma(t))$ {forward diffusion}
4:  $\mathcal{L}_{\text{guidance}}\!=\!\dfrac{1}{b}\sum\|r_{\phi}(x_{t},t)-\dfrac{1-d_{\omega}(x_{0})}{d_{\omega}(x_{0})}\|_{2}^{2}$ {Guidance loss}
5:
6:  Sample mini‑batch $x_{0}$ of size $b$ from $\mathbf{q}$ and draw $t\sim\text{Uniform}(0,1)$
7:  $x_{t}\leftarrow\textsc{Corrupt}(x_{0},\sigma(t))$ {forward diffusion}
8:  $\mathcal{L}_{\text{cycle}}\!=\!\dfrac{1}{b}\sum\|r_{\phi}(x_{t},t)-\dfrac{1-d_{\omega}(x_{t},t)}{d_{\omega}(x_{t},t)}\|_{2}^{2}$ {Guidance loss}
9:
10:  $\mathcal{L}=\mathcal{L}_{\text{guidance}}+\lambda\cdot\mathcal{L}_{\text{cycle}}$
11:  $\psi\leftarrow\psi-\eta\,\nabla_{\psi}\mathcal{L}$ {gradient update}
12:until convergence
13:
14:return trained parameters $\psi$
For sample corruption, we use a log-linear schedule with $\sigma_{\text{max}}=20$ and $\sigma_{\text{min}}=0.0001$. The planner is trained following Algorithm 4, using the same training configuration as described above, except that it is trained for 100k update steps. On the held-out validation set from the source domain, the planner achieves an average accuracy of 70%. In high-noise regions, it becomes considerably more challenging for the planner to predict which positions will be correctly denoised, due to the high variability in token unmasking.
Algorithm 4 Training loop for the planner $\rho_{\vartheta}(x_{t},t)$
[…]
1:repeat
2:  Sample mini-batch $x_{0}$ of size $b$ from $\mathcal{D}$; draw $t\sim\mathrm{Uniform}(0,1)$
3:  $x_{t}\leftarrow\textsc{Corrupt}(x_{0},\sigma(t))$ {(1) corrupt the sample}
4:  $\log x_{\theta}\leftarrow x_{\theta}(x_{t})$;   $\hat{x}_{0}\leftarrow\arg\max\log x_{\theta}$ {(2) pass to source denoiser}
5:  $M\leftarrow(x_{t}=\texttt{[MASK]})$;   $y\leftarrow\mathbf{1}\!\left[\hat{x}_{0}=x_{0}\right]$ {(3) where denoiser is correct (masked positions)}
6:  $s\leftarrow\rho_{\vartheta}(x_{t})$ {(4) planner logits per position}
7:  $\mathcal{L}\leftarrow\mathrm{BCEWithLogits}\!\left(s[M],\,y[M]\right)$ {(5) loss over masked positions only}
8:  $\vartheta\leftarrow\vartheta-\eta\nabla_{\vartheta}\mathcal{L}$ {gradient update}
9:until convergence
10:
11:return trained planner parameters $\vartheta$
Appendix D Application to continuous-time Score-based Models
D.1 Background
(campbell2022continuoustimeframeworkdiscrete) introduces a continuous-time framework for discrete denoising models formulated as a continuous-time Markov chain (CTMC). Unlike the discrete-time setting of (austin2023structureddenoisingdiffusionmodels) with steps $t=0,1,\dots,T$, here $t\in[0,T]$ is continuous.
In this section, we switch notation. The forward kernel is $p_{t|s}(\mathbf{y}^{\prime}\mid\mathbf{y}):=p(\mathbf{z}_{t}=\mathbf{y}^{\prime}\mid\mathbf{z}_{s}=\mathbf{y})$, the backward kernel is $p_{s|t}(\mathbf{y}^{\prime}\mid\mathbf{y}):=p(\mathbf{z}_{s}=\mathbf{y}^{\prime}\mid\mathbf{z}_{t}=\mathbf{y})$, the noising kernel is $p_{t}(\mathbf{y}\mid\mathbf{x}):=p(\mathbf{z}_{t}=\mathbf{y}\mid\mathbf{x}_{0}=\mathbf{x})$, and the true backward transition is $p(\mathbf{y}^{\prime}\mid\mathbf{y},\mathbf{x}):=p(\mathbf{z}_{s}=\mathbf{y}^{\prime}\mid\mathbf{z}_{t}=\mathbf{y},\mathbf{x}_{0}=\mathbf{x})$. This CTMC formulation allows transitions at arbitrary times. The forward corruption is specified by a rate matrix $R_{t}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, and the reverse process by a reverse rate matrix $\widetilde{R}_{t}$:
$p_{t|s}(\mathbf{y}^{\prime}\mid\mathbf{y})$ $=\delta_{\mathbf{y}^{\prime},\mathbf{y}}+R_{t}(\mathbf{y}^{\prime},\mathbf{y})\,\Delta t+\mathcal{O}\!\left(\Delta t^{2}\right),$
$p_{s|t}(\mathbf{y}^{\prime}\mid\mathbf{y},\mathbf{x})$ $=\delta_{\mathbf{y}^{\prime},\mathbf{y}}+\widetilde{R}_{t}(\mathbf{y}^{\prime},\mathbf{y})\,\Delta t+\mathcal{O}\!\left(\Delta t^{2}\right).$
Here $\Delta t$ denotes a small time increment when discretizing the continuous-time process, with $s=t-\Delta t$. Here, $\delta_{i,j}$ denotes the Kronecker delta (equal to $1$ if $i=j$ and 0 otherwise), and $\mathcal{O}(\Delta t^{2})$ denotes terms of order $\Delta t^{2}$ or higher. The reverse rate can be written in terms of the forward rate and the marginals $q_{t}(\cdot\mid\mathbf{x})$:
$\widetilde{R}_{t}(\mathbf{y}^{\prime},\mathbf{y})=\begin{cases}\dfrac{q_{t}(\mathbf{y}^{\prime}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}\,R_{t}(\mathbf{y},\mathbf{y}^{\prime}),&\mathbf{y}^{\prime}\neq\mathbf{y},\\[6.0pt] -\displaystyle\sum_{\tilde{\mathbf{y}}\neq\mathbf{y}}\dfrac{q_{t}(\tilde{\mathbf{y}}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}\,R_{t}(\mathbf{y},\tilde{\mathbf{y}}),&\mathbf{y}^{\prime}=\mathbf{y}.\end{cases}$
Concrete score.
We can express the reverse process in terms of the forward transition rates and a concrete score defined by $s_{\theta}(\mathbf{y})_{\mathbf{y}^{\prime}}\approx\tfrac{q_{t}(\mathbf{y}^{\prime}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}$, parameterized by a neural network $s_{\theta}:\mathcal{V}\to\mathbb{R}_{>0}^{|\mathcal{V}|}$. This enables training via concrete score matching (CSM) (meng2023concretescorematchinggeneralized), directly analogous to score-based models in continuous domains (song2021scorebasedgenerativemodelingstochastic). (lou2024discretediffusionmodelingestimating) propose the diffusion-weighted denoising score entropy (DWDSE) loss (their CDM-based objective) and note it can also be derived from the continuous-time likelihood framework of (campbell2022continuoustimeframeworkdiscrete); a detailed derivation appears in Appendix C.3 of (sahoo2024simpleeffectivemaskeddiffusion). Concretely,
$\lim_{T\to\infty}\mathbb{E}_{\begin{subarray}{c}t\in\{\tfrac{1}{T},\tfrac{2}{T},\ldots,1\}\\ \mathbf{y}\sim q_{t}(\cdot\mid\mathbf{x})\end{subarray}}\!\bigl[\mathrm{D}_{\mathrm{KL}}\!\bigl(p_{s\mid t}(\mathbf{y}^{\prime}\mid\mathbf{y},\mathbf{x})\;\|\;p_{s\mid t}(\mathbf{y}^{\prime}\mid\mathbf{y})\bigr)\bigr]$ (36)
$=\mathbb{E}_{\begin{subarray}{c}t\in[0,1]\\ \mathbf{y}\sim q_{t}(\cdot\mid\mathbf{x})\end{subarray}}\!\Bigl[\sum_{\mathbf{y}^{\prime}\neq\mathbf{y}}R_{t}(\mathbf{y},\mathbf{y}^{\prime})\Bigl(s_{\theta}(\mathbf{y})_{\mathbf{y}^{\prime}}-\frac{q_{t}(\mathbf{y}^{\prime}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}\log s_{\theta}(\mathbf{y})_{\mathbf{y}^{\prime}}+K\!\Bigl(\tfrac{q_{t}(\mathbf{y}^{\prime}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}\Bigr)\Bigr)\Bigr],$ (37)
where $K(a)=a\log a-a$.
D.2 Theorem
In the previous section, we introduced the diffusion process and the source-domain objective on which the score function $s_{\theta}$ is trained. We now state a theorem that enables sampling from the target distribution without retraining $s_{\theta}$ under the objective in 37. Let $r_{\phi}(\mathbf{y})$ denote the learned density-ratio guidance term defined earlier.
Theorem 3 (Score-Based Ratio Transfer Learning)
Let $s_{\theta}(\mathbf{y})_{\mathbf{y}^{\prime}}$ be a score-based model trained on the source distribution $p$, and let $\widetilde{R}_{t}^{\theta}(\cdot,\cdot)$ be the reverse-rate matrix induced by $s_{\theta}$. Then, sampling from the target distribution $q$ is given by the target reverse-rate matrix:
$\widetilde{R}_{t}^{\psi}(\mathbf{y}^{\prime},\mathbf{y})=\begin{cases}\dfrac{r_{\phi}(\mathbf{y}^{\prime})}{r_{\phi}(\mathbf{y})}\,\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y}),&\mathbf{y}^{\prime}\neq\mathbf{y},\\[8.0pt] -\displaystyle\sum_{\tilde{\mathbf{y}}\neq\mathbf{y}}\dfrac{r_{\phi}(\tilde{\mathbf{y}})}{r_{\phi}(\mathbf{y})}\,\widetilde{R}_{t}^{\theta}(\tilde{\mathbf{y}},\mathbf{y}),&\mathbf{y}^{\prime}=\mathbf{y}.\end{cases}$ (38)
D.3 Proof of Theorem 3
For the proof, we take the loss for a small step $\Delta t$ with $s=t-\Delta t$ from (36) and apply Theorem 1, which yields the following expression:
$q_{s\mid t}^{\psi^{\star}}(\mathbf{y}^{\prime}\mid\mathbf{y})$ $=\frac{p_{s\mid t}(\mathbf{y}^{\prime}\mid\mathbf{y})\,\mathbb{E}_{\mathbf{x}\sim p(\cdot\mid\mathbf{y}^{\prime})}\!\bigl[\tfrac{q(\mathbf{x})}{p(\mathbf{x})}\bigr]}{\displaystyle\sum_{\widetilde{\mathbf{y}}}p_{s\mid t}(\widetilde{\mathbf{y}}\mid\mathbf{y})\,\mathbb{E}_{\mathbf{x}\sim p(\cdot\mid\widetilde{\mathbf{y}})}\!\bigl[\tfrac{q(\mathbf{x})}{p(\mathbf{x})}\bigr]}$ (39)
$=\frac{p_{s\mid t}(\mathbf{y}^{\prime}\mid\mathbf{y})\,r_{\phi}(\mathbf{y}^{\prime})}{\displaystyle\sum_{\widetilde{\mathbf{y}}}p_{s\mid t}(\widetilde{\mathbf{y}}\mid\mathbf{y})\,r_{\phi}(\widetilde{\mathbf{y}})}$ (40)
We can already assume that the source backward kernel is approximated by $p_{s\mid t}(\mathbf{y}^{\prime}\mid\mathbf{y})=\delta_{\mathbf{y}^{\prime},\mathbf{y}}+\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y})\,\Delta t+\mathcal{O}(\Delta t^{2})$:
$q_{s\mid t}^{\psi}(\mathbf{y}^{\prime}\mid\mathbf{y})=\frac{\bigl[\delta_{\mathbf{y}^{\prime},\mathbf{y}}+\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y})\Delta t\bigr]\,r_{\phi}(\mathbf{y}^{\prime})}{\sum_{\tilde{\mathbf{y}}}\bigl[\delta_{\tilde{\mathbf{y}},\mathbf{y}}+\widetilde{R}_{t}^{\theta}(\tilde{\mathbf{y}},\mathbf{y})\Delta t\bigr]\,r_{\phi}(\tilde{\mathbf{y}})}\;+\;\mathcal{O}(\Delta t^{2})$
[…]
$\frac{1}{\sum_{\tilde{\mathbf{y}}}\bigl[\delta_{\tilde{\mathbf{y}},\mathbf{y}}+\widetilde{R}_{t}^{\theta}(\tilde{\mathbf{y}},\mathbf{y})\Delta t\bigr]\,r_{\phi}(\tilde{\mathbf{y}})}=\frac{1}{r_{\phi}(\mathbf{y})+\Delta t\,S}=\frac{1}{r_{\phi}(\mathbf{y})}\Bigl(1+\frac{\Delta t\,S}{r_{\phi}(\mathbf{y})}\Bigr)^{-1}=\frac{1}{r_{\phi}(\mathbf{y})}-\frac{\Delta t\,S}{r_{\phi}(\mathbf{y})^{2}}+\mathcal{O}(\Delta t^{2})$
  1. 1.
    For the Off–diagonal case $\mathbf{y}^{\prime}\neq\mathbf{y}$, hence $\delta_{\mathbf{y}^{\prime},\mathbf{y}}=0$:
    $q_{s\mid t}^{\psi}(\mathbf{y}^{\prime}\mid\mathbf{y})$ $=\frac{\bigl[\delta_{\mathbf{y}^{\prime},\mathbf{y}}+\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y})\Delta t\bigr]\,r_{\phi}(\mathbf{y}^{\prime})}{\sum_{\tilde{\mathbf{y}}}\bigl[\delta_{\tilde{\mathbf{y}},\mathbf{y}}+\widetilde{R}_{t}^{\theta}(\tilde{\mathbf{y}},\mathbf{y})\Delta t\bigr]\,r_{\phi}(\tilde{\mathbf{y}})}\;+\;\mathcal{O}(\Delta t^{2})$
    $=\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y})\,r_{\phi}(\mathbf{y}^{\prime})\,\Delta t\left(\frac{1}{r_{\phi}(\mathbf{y})}-\frac{\Delta t\,S}{r_{\phi}(\mathbf{y})^{2}}+\mathcal{O}(\Delta t^{2})\right)$
    $=\frac{\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y})\,r_{\phi}(\mathbf{y}^{\prime})}{r_{\phi}(\mathbf{y})}\,\Delta t\;+\;\mathcal{O}(\Delta t^{2}),$
  2. 2.
    For the diagonal case $\mathbf{y}^{\prime}=\mathbf{y}$:
    $q_{s\mid t}^{\psi}(\mathbf{y}\mid\mathbf{y})$ $=\frac{\bigl[1+\widetilde{R}_{t}^{\theta}(\mathbf{y},\mathbf{y})\,\Delta t\bigr]\,r_{\phi}(\mathbf{y})}{r_{\phi}(\mathbf{y})+\Delta t\,S}+\mathcal{O}(\Delta t^{2})=\frac{1+\widetilde{R}_{t}^{\theta}(\mathbf{y},\mathbf{y})\,\Delta t}{1+\frac{\Delta t\,S}{r_{\phi}(\mathbf{y})}}+\mathcal{O}(\Delta t^{2})$
    $=\Bigl(1+\widetilde{R}_{t}^{\theta}(\mathbf{y},\mathbf{y})\,\Delta t\Bigr)\Bigl(1-\frac{\Delta t\,S}{r_{\phi}(\mathbf{y})}\Bigr)+\mathcal{O}(\Delta t^{2})=1+\Bigl[\widetilde{R}_{t}^{\theta}(\mathbf{y},\mathbf{y})-\frac{S}{r_{\phi}(\mathbf{y})}\Bigr]\Delta t+\mathcal{O}(\Delta t^{2}).$
Putting these two cases together, we arrive at the following:
$q_{s\mid t}^{\psi}(\mathbf{y}^{\prime}\mid\mathbf{y})$ $=\delta_{\mathbf{y}^{\prime},\mathbf{y}}\Bigl[1+\bigl(\widetilde{R}_{t}^{\theta}(\mathbf{y},\mathbf{y})-\sum_{\tilde{\mathbf{y}}}\widetilde{R}_{t}^{\theta}(\tilde{\mathbf{y}},\mathbf{y})\,\tfrac{r_{\phi}(\tilde{\mathbf{y}})}{r_{\phi}(\mathbf{y})}\bigr)\Delta t\Bigr]+(1-\delta_{\mathbf{y}^{\prime},\mathbf{y}})\Bigl[\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y})\,\tfrac{r_{\phi}(\mathbf{y}^{\prime})}{r_{\phi}(\mathbf{y})}\,\Delta t\Bigr]+\mathcal{O}(\Delta t^{2})$
$=\delta_{\mathbf{y}^{\prime},\mathbf{y}}+{\Delta t}\Bigl[\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y})\,\frac{r_{\phi}(\mathbf{y}^{\prime})}{r_{\phi}(\mathbf{y})}-\delta_{\mathbf{y}^{\prime},\mathbf{y}}\,\sum_{\tilde{\mathbf{y}}}{\widetilde{R}_{t}^{\theta}(\tilde{\mathbf{y}},\mathbf{y})\,\frac{r_{\phi}(\tilde{\mathbf{y}})}{r_{\phi}(\mathbf{y})}}\Bigr]+\mathcal{O}(\Delta t^{2}),$
$=\delta_{\mathbf{y}^{\prime},\mathbf{y}}+\Delta t\,\widetilde{R}_{t}^{\psi}(\mathbf{y}’,\mathbf{y})\,+\mathcal{O}(\Delta t^{2}).$
Thus concluding the proof.
[…]
The work from (nisonoff2025unlockingguidancediscretestatespace) in equation (2) arrives at a similar term just for classifier-based guidance.

### v2

@@
 Abstract
-Discrete diffusion models achieve strong performance across language and other discrete domains, providing a powerful alternative to autoregressive models.
-However, their strong performance relies on large training datasets, which are costly or risky to obtain, especially when adapting to new domains.
-Transfer learning is the natural way to adapt pretrained discrete diffusion models, but current methods require fine-tuning large diffusion models, which is computationally expensive and often impractical.
-Building on ratio-based transfer learning for continuous diffusion, we provide Guided Transfer Learning for discrete diffusion models (GTL).
-This enables sampling from a target distribution without modifying the pretrained denoiser.
-The same guidance formulation applies to both discrete-time diffusion and continuous-time score-based discrete diffusion, yielding a unified treatment.
-Guided discrete diffusion often requires many forward passes of the guidance network, which becomes impractical for large vocabularies and long sequences.
-To address this, we further present an efficient guided sampler that concentrates evaluations on planner-selected positions and top candidate tokens, thus lowering sampling time and computation.
-This makes guided language modeling practical at scale for large vocabularies and long sequences.
-We evaluate GTL on sequential data, including synthetic Markov chains and language modeling, and provide empirical analyses of its behavior.
-This is a pre-print.
+Discrete diffusion models (DMs) have achieved strong performance in language and other discrete domains, offering a compelling alternative to autoregressive modeling.
+Yet this performance typically depends on large training datasets, challenging the performance of DMs in small-data regimes—common under real-world constraints.
+Aimed at this challenge, recent work in continuous DMs suggests that transfer learning via classifier ratio–based guidance can adapt a pretrained DM to a related target distribution, often outperforming alternatives such as full-weight fine-tuning on the target data.
+By contrast, transfer learning for discrete DMs remains unexplored.
+We address this gap by exploring practical analogues of ratio-based transfer learning for discrete DMs.
+Our theoretical analysis shows that a direct extension of existing ratio-based guidance is computationally prohibitive, scaling with vocabulary size.
+To overcome this limitation, we introduce a scheduling mechanism that yields a practical algorithm, Guided Transfer Learning for discrete diffusion models (GTL).
+GTL enables sampling from a target distribution without modifying the pretrained denoiser and reduces the cost to linear scaling in vocabulary size, which in turn supports longer sequence generation.
+We evaluate GTL on sequential data, including synthetic Markov chains and language modeling tasks, and provide a detailed empirical analysis of its behavior.
+The results highlight a clear trade-off: when target datasets are large, weight fine-tuning is often preferable, whereas GTL becomes increasingly effective as target data shrinks.
+Finally, we experimentally demonstrate a key failure mode of GTL: when the source and target distributions overlap poorly, the ratio-based classifier required for guidance becomes unreliable, limiting transfer performance.
 1 Introduction
-Diffusion models (ho2020denoisingdiffusionprobabilisticmodels; song2021scorebasedgenerativemodelingstochastic) have become the dominant approach for generative modeling in continuous domains like image, video, and audio (lovelace2023latentdiffusionlanguagegeneration; ho2022videodiffusionmodels; liu2023audioldmtexttoaudiogenerationlatent).
-More recently, discrete diffusion models for language have achieved competitive performance and, by denoising tokens in parallel and in any order, offer practical advantages for controllable infilling and editing compared to left-to-right autoregressive decoding (gong2025scalingdiffusionlanguagemodels; li2022diffusionlm).
-Among existing discrete diffusion approaches, masked diffusion models (MDMs) have demonstrated remarkable performance in tasks such as reasoning (nie2025scalingmaskeddiffusionmodels; zheng2025maskeddiffusionmodelssecretly), planning (ye2025autoregressiondiscretediffusioncomplex), and infilling (gong2025scalingdiffusionlanguagemodels), often performing on par with state-of-the-art auto-regressive language models (ALMs).
-An open problem in the field is to develop transfer learning approaches that enable leveraging the incredible performance of large pre-trained diffusion models in domains with scarce data (ouyang2024transferlearningdiffusionmodels).
-In the continuous domain, recent works have explored transfer learning approaches based domain classifiers (ratio estimators) (zhong2025domainguidancesimpletransfer; ouyang2024transferlearningdiffusionmodels).
-Such methods bypass the need of updating millions of parameters, providing an alternative to adapter-based finetuning (moon2022finetuning; xie2023difffitunlockingtransferabilitylarge) and reinforcement learning (fan2023dpokreinforcementlearningfinetuning; clark2024directlyfinetuningdiffusionmodels) in scenarios of extreme scarcity.
-In this paper, we tackle the problem of extending domain classifier-based sampling for discrete diffusion models.
-We first develop the theoretical analogues of (ouyang2024transferlearningdiffusionmodels), leading to a first naïve implementation of (8) (In supplementary materials, algorithm 2).
-However, we show that such naïve implementation requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per denoising step, scaling poorly with sequence length $L$ and vocabulary size $|\mathcal{V}|$, limiting its applicability to domains such as language modeling.
-Motivated by these challenges, we introduce Guided Transfer Learning (GTL) for discrete diffusion models.
-Our method leverages a pretrained source discrete diffusion model and introduces a compact ratio network that estimates the density ratio between the target and source data distributions.
-We demonstrate that sampling from the target domain can be achieved solely by training this ratio-based guidance module, enabling effective target sampling without updating the denoiser.
-This substantially reduces the training cost compared to conventional fine-tuning approaches.
-We establish a closed-form solution for the sampling method, and we further extend this argument to continuous-time score-based models.
-This generalization allows our approach to cover the two dominant diffusion frameworks: score-based models (lou2024discretediffusionmodelingestimating; sun2023scorebasedcontinuoustimediscretediffusion) and discrete-time diffusion models (austin2023structureddenoisingdiffusionmodels), making the method broadly applicable to diffusion modeling.
-This GTL construction does not follow directly from continuous diffusion models (ouyang2024transferlearningdiffusionmodels), since score-matching diffusion is not defined for categorical discrete tokens and cannot be derived from the score-matching loss function (song2021scorebasedgenerativemodelingstochastic).
-We further introduce a guided sampling algorithm that reduces the number of guidance calls per denoising step by focusing computation on planner-selected tokens.
-This design improves efficiency.
-Previewing our results, GTL surpasses vanilla and fine-tuned diffusion across all data-scarcity regimes while training only  7% as many parameters (Fig.
+Diffusion models (DMs) (Ho et al., 2020; Song et al., 2021) have become the dominant approach for generative modeling in continuous domains like image, video, and audio (Lovelace et al., 2023; Ho et al., 2022; Liu et al., 2023).
+More recently, discrete DMs for language have achieved competitive performance and, by denoising tokens in parallel and in any order, offer practical advantages for controllable infilling and editing compared to left-to-right autoregressive decoding (Gong et al., 2025; Li et al., 2022a).
+Among existing discrete diffusion approaches, masked diffusion models (MDMs) have demonstrated remarkable performance in tasks such as reasoning (Nie et al., 2025; Zheng et al., 2025), planning (Ye et al., 2025), and infilling (Gong et al., 2025), often performing on par with state-of-the-art auto-regressive language models (ALMs).
+Training DMs require large amounts of data (Wang et al., 2023).
+And it remains an open question how to best leverage their powerful generative modeling capabilities in domains where scarse data is available (Zhang et al., 2025), a common scenario in applications like medical imaging, where privacy and limited public availability constrain dataset size (Xie et al., 2025).
+This problem motivates the exploration of transfer learning techniques, where the goal is leverage a DM pretrained on a large dataset for generative modeling in a related domain where data is more scarse.
+One possible solution is to fine-tune the pretrained DM’s denoiser on the target data (Xie et al., 2023b), but this becomes ineffective in very small data regimes (Moon et al., 2022).
+To address this, recent works have proposed transfer learning to tackle the small data regime by using classifier-based guidance (CBG) (Zhong et al., 2025; Ouyang et al., 2024), which estimates a density-ratio between the source and target domain and uses it as a diffusion guidance mechanism.
+Such methods are more effective than finetuning in scenarios of extreme scarcity, as they bypass the need of updating millions of parameters (Moon et al., 2022; Xie et al., 2023a).
+However, these methods have only been investigated in continuous domains (e.g., images).
+By contrast, transfer learning for discrete DMs such as language models, remains largely unexplored.
+Outside of transfer learning, Schiff et al.
+(2025); Nisonoff et al.
+(2025) have explored CBG mechanisms for discrete DMs.
+These methods apply mainly to domains with small $|\mathcal{V}|$ and modest $L$, such as biological sequences, leaving scalable CBG for language modeling as an open problem.
+This constaints are due to the necessity of token substitutions across all positions and vocabulary items during score computation.
+More precisely, letting $L$ be the sequence length and $\mathcal{V}$ the vocabulary, naive CBG requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies.
+In this paper, we aim to leverage CBGs for transfer learning, while overcoming the limiting small vocabulary sizes.
+First, we present a theoretical analysis deriving a closed-form ratio-guided reverse transition for discrete-time diffusion and extend this principle to continuous-time, score-based discrete diffusion, covering both score-based models (Lou et al., 2024; Sun et al., 2023) and discrete-time diffusion models (Austin et al., 2023).
+Such derivation is architecture- and corruption-agnostic—deriving the result from the evidence lower bound (ELBO).
+Next, we address the scaling problem of CBGs for discrete DMs.
+For this purpose, we introduce an efficient sampler that makes guided language modeling for masked discrete DMs practical by avoiding $\mathcal{O}(L|\mathcal{V}|)$ guidance calls, concentrating ratio evaluations on planner-selected positions and a top-$n_{\text{ratio}}$ candidate set.
+[…]
+Leveraging this result, we introduce a method termed Guided Transfer Learning (GTL).
+The core idea of GTL is to freeze a pretrained source denoiser and learn a lightweight ratio network from mixed source/target data to estimate target-to-source density ratios; at sampling time, we use this ratio with a learned planner network as guidance to reweight the denoiser’s reverse transition and generate target-domain samples without fine-tuning.
+This procedure allows to effectively utilize transfer learning based on CBG for discrete DMs.
+Previewing our results, GTL outperforms vanilla and fine-tuned diffusion across data-scarcity regimes while training only $\sim 7\%$ as many parameters (Fig.
 1).
+In a setting common in practice—where the pretraining corpus already contains some in-domain target samples—we find that GTL performs better and yields more stable performance.
+Our result also reveals limitations of CBG for discrete modeling: under weak domain overlap, the interaction between guidance strength $\gamma$ and candidate pruning can amplify ratio errors.
+In this regime, larger $\gamma$ or small $n_{\text{ratio}}$ may destabilize denoising and hurt quality.
 We summarize our contributions as follows:
@@
     We validate GTL and the guided sampler on sequential data, including synthetic Markov chains and arXiv-abstract language modeling tasks.
-Figure 1: GTL outperforms with  7% of the parameters.
-MAUVE ($\uparrow$) vs.
-fraction of target-domain training data (e.g., 100% = all 79,631 arXiv Physics abstracts; $1\%\approx 800$).
-Our Guided Transfer Learning (GTL, green) achieves the highest MAUVE across all data regimes while training only a 4.1M parameter ratio model outperforming both Vanilla (trained only on target data) and Fine-tuned (source-pretrained then fine-tuned on target) models.
-Points show mean, and error bars show standard error over 3×3 (train×sample) seeds.
-See Section 5.2 for the experimental setup and evaluation details.
-2 Related Work
-Discrete Diffusion Models
-Discrete diffusion models (DDMs) are widely used for biological sequence generation (e.g., DNA, RNA, proteins) as well as language modeling (Sarkar2024; yu2025discretediffusionlargelanguage).
-Continuous‑noise variants that corrupt embeddings or latents are likely to suffer from embedding‑to‑token projection mismatch  (li2022diffusionlmimprovescontrollabletext; lovelace2023latentdiffusionlanguagegeneration).
-Beyond that, two widely adopted formulations operate entirely in discrete space, avoiding continuous mappings.
-First, simplified masked diffusion models such as MLDM and MD4 follow the D3PM framework, using discrete‑time transition matrices with a denoiser trained to invert the forward kernels (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion; austin2023structureddenoisingdiffusionmodels).
-Second, continuous‑time discrete diffusion estimates a discrete score, which represents the log‑ratio between adjacent‑timestep marginals.
-This induces a continuous-time Markov chain (CTMC) (lou2024discretediffusionmodelingestimating; meng2023concretescorematchinggeneralized; campbell2022continuoustimeframeworkdiscrete).
-Guided Diffusion Models
-In both continuous and discrete settings, guidance strategies like classifier‑based or classifier‑free are widely used.
-In the discrete case, classifier‑free guidance typically performs two forward passes at inference time, one conditional and one unconditional, while training a single large model on a mixture of conditional prompt–input pairs and unconditional data (huang2025ctrldiffboostinglargediffusion).
-Closer to our approach is classifier‑based guidance, which steers generation using a classifier $p(y\mid x_{t},t)$ while the denoiser is trained on the full corpus; sampling is biased toward the target label $y$ in the data distribution (schiff2025simpleguidancemechanismsdiscrete; nisonoff2025unlockingguidancediscretestatespace).
-In both discrete- and continuous-time formulations, this typically yields per-token factorized guidance, which incurs many computational steps, scaling with $\mathcal{O}(L|\mathcal{V}|)$ evaluations per denoising step.
-To mitigate this cost, (nisonoff2025unlockingguidancediscretestatespace) proposes a first-order Taylor approximation that reduces guidance evaluations to a forward and backward pass.
-TESS2 (tae2025tess2largescalegeneralist) instead applies gradient ascent on a reward with respect to the model logits during inference to improve response quality.
-Finally, the Energy‑based Diffusion Language Model (xu2025energybaseddiffusionlanguagemodels) defines an energy over predicted clean sequences and uses it to importance‑weight the denoising step, reducing the train–sample mismatch in discrete diffusion models.
-Transfer Learning
-Fine-tuning and transfer learning share the goal of adapting a model to a new task or domain.
-Fine-tuning is computationally expensive because it updates the pretrained denoiser’s weights on new data.
-Prior work on continuous diffusion models explores reinforcement learning (fan2023dpokreinforcementlearningfinetuning; clark2024directlyfinetuningdiffusionmodels) and adapter-based fine-tuning (moon2022finetuning; xie2023difffitunlockingtransferabilitylarge).
-In contrast, transfer learning for diffusion remains comparatively underexplored.
-Transfer Learning for Diffusion Models (TLDM) proposes a ratio-estimation technique for continuous score-based models, but demonstrates effectiveness on synthetic data and a one-dimensional heartbeat dataset, leaving applications to higher-dimensional and large-scale domains open.
 2 Background
@@
 Discrete diffusion
-Discrete Denoising Diffusion Probabilistic Models (D3PMs) (austin2023structureddenoisingdiffusionmodels) provide a flexible framework for diffusion over discrete state spaces.
+Discrete Denoising Diffusion Probabilistic Models (D3PMs) (Austin et al., 2023) provide a flexible framework for diffusion over discrete state spaces.
 The forward Markov chain is $p(\mathbf{z}_{t}\mid\mathbf{z}_{s})=\text{Cat}\!\left(\mathbf{z}_{t};\,Q_{t\mid s}\mathbf{z}_{s}\right)$ where $Q_{t\mid s}\in\mathbb{R}^{N\times N}$ is a transition matrix whose $(i,j)$ entry is the probability of moving to state $i$ at time $t$ from state $j$ at time $s$.
@@
 Masked (absorbing-state) diffusion.
-The absorbing-state variant or often called masked diffusion, replaces tokens with a special [MASK] token (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion).
+The absorbing-state variant or often called masked diffusion, replaces tokens with a special [MASK] token (Sahoo et al., 2024; Shi et al., 2025).
 It can be viewed as an interpolation between the clean distribution and a noise prior over the vocabulary.
@@
 Making use of the masking process, the backward posterior simplifies to:
-$p(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})$
-$=\begin{cases}\operatorname{Cat}\!\left(\mathbf{z}_{s};\,\mathbf{z}_{t}\right),&\text{if }\mathbf{z}_{t}\neq\mathbf{m},\\[6.0pt] \operatorname{Cat}\!\left(\mathbf{z}_{s};\,\dfrac{(1-\alpha_{s})\mathbf{m}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{0}}{1-\alpha_{t}}\right),&\text{if }\mathbf{z}_{t}=\mathbf{m}.\end{cases}$ (2)
+$p(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})=\begin{cases}\operatorname{Cat}\!\left(\mathbf{z}_{s};\,\mathbf{z}_{t}\right),&\text{if }\mathbf{z}_{t}\neq\mathbf{m},\\[6.0pt] \operatorname{Cat}\!\left(\mathbf{z}_{s};\,\dfrac{(1-\alpha_{s})\mathbf{m}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{0}}{1-\alpha_{t}}\right),&\text{if }\mathbf{z}_{t}=\mathbf{m}.\end{cases}$ (2)
 Hence, when a token is unmasked ($\mathbf{x}_{t}\neq\mathbf{m}$) the step is deterministic: $\mathbf{x}_{s}=\mathbf{x}_{t}$.
 Otherwise, the posterior is a linear interpolation between the mask distribution and the clean token.
-This structure yields the simplified diffusion loss (sahoo2024simpleeffectivemaskeddiffusion; shi2025simplifiedgeneralizedmaskeddiffusion):
-$\mathcal{L}_{\text{diffusion}}$ $=\sum_{i=1}^{T}D_{\mathrm{KL}}\!\left(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\right)$ (3)
-$=\sum_{i=1}^{T}\mathbb{E}_{q}\!\left[\frac{\alpha_{t(i)}-\alpha_{s(i)}}{1-\alpha_{t(i)}}\log\big\langle\mathbf{x}_{\theta}(\mathbf{z}_{t(i)}),\mathbf{x}_{0}\big\rangle\right],$ (4)
+This structure yields the simplified diffusion loss (Sahoo et al., 2024; Shi et al., 2025):
+$\mathcal{L}_{\text{diffusion}}$ $=\sum_{i=1}^{T}D_{\mathrm{KL}}\!\left(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\big\|\,p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\right)=\sum_{i=1}^{T}\mathbb{E}_{q}\!\left[\frac{\alpha_{t(i)}-\alpha_{s(i)}}{1-\alpha_{t(i)}}\log\big\langle\mathbf{x}_{\theta}(\mathbf{z}_{t(i)}),\mathbf{x}_{0}\big\rangle\right],$ (3)
 where $\mathbf{x}$ is the one-hot encoding of the clean data and $\langle\cdot,\cdot\rangle$ is the inner product.
@@
 In this work, we focus on the classifier-based variant, which allows us to rearrange the conditional generative process using Bayes’ rule as $p^{\gamma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},y)=p(y\mid\mathbf{z}_{s},\mathbf{z}_{t})^{\gamma}\,p(\mathbf{z}_{s}\mid\mathbf{z}_{t})/Z$, with $Z$ being a normalizing constant.
+(Schiff et al., 2025; Nisonoff et al., 2025)
 For the extension to sequences of tokens $\mathbf{z}_{s}^{(1:L)}$ of length $L$, we assume that the distribution $p^{\gamma}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)},y\right)$ factorizes independently across tokens.
 This leads to the following formulation:
-$p^{\gamma}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)},y\right)$
-$=\prod_{\ell=1}^{L}\frac{p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}p\!\left(y\mid\tilde{\mathbf{z}}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}$ (5)
+$p^{\gamma}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)},y\right)=\prod_{\ell=1}^{L}\frac{p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}p\!\left(y\mid\tilde{\mathbf{z}}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)^{\gamma}\,p\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}$ (4)
 Here, $\mathcal{V}$ denotes the vocabulary set, and $p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)$ represents the classifier probability given the sequence $\mathbf{z}_{t}^{(1:L)}$ with the token $\mathbf{z}_{s}$ placed at position $\ell$.
-However, this formulation comes with the drawback of requiring $\mathcal{O}(L\cdot|\mathcal{V}|)$ forward passes through the classifier model.
+However, this formulation comes with the drawback of requiring $\mathcal{O}(L\cdot|\mathcal{V}|)$ forward passes through the classifier model, which is infeasible for long sequences and large vocabularies.
+Because each evaluation corresponds to a distinct single-token intervention on the full sequence, this cost is not removed by straightforward parallelization.
 Problem Formulation
@@
 3 Guided Transfer Learning for DDMs
-We propose Guided Transfer Learning for Discrete Diffusion Models, a framework that enables transfer learning for discrete diffusion models.
-We first formalize the problem setup, then present the mathematical formulation of our method, and finally introduce a new sampling algorithm that leverages our results to make transfer learning broadly applicable across datasets.
-Given the problem statement and access to both domains, one could fine-tune the source network $p_{\theta}$ by maximizing (4), thereby optimizing the entire parameter set.
-For large language models, this is often costly and slow due to the scale of parameters.
-Moreover, when the target domain provides far fewer samples, fine-tuning alone rarely yields a strong generative model in terms of diversity and quality.
-Our method is based on the following theorem, which proposes a different approach to this problem: it leverages a pretrained source-domain diffusion model together with a ratio estimator between the two domains.
-We assume the two diffusion models share the same forward process $p(\mathbf{z}_{t}\mid\mathbf{x}_{0})=q(\mathbf{z}_{t}\mid\mathbf{x}_{0})$ for all $t$, that the density ratio $q(\mathbf{x}_{0})/p(\mathbf{x}_{0})$ is well defined.
-Relation to TLDM.
-TLDM (ouyang2024transferlearningdiffusionmodels) proves a transfer result for continuous, score-based diffusion and is not directly applicable to discrete state spaces.
-The theorem is stated in $\mathbb{R}^{d}$ and relies on score matching with differentiable densities, requiring gradients such as $\nabla_{x_{t}}\log q_{t}(x_{t})$.
-In the discrete setting (categorical tokens), such Euclidean scores are undefined.
-Beyond this, the modeling and training objectives differ: continuous diffusion injects Gaussian noise and optimizes score-matching Mean Squared Error, whereas discrete diffusion uses categorical transition matrices (e.g., absorbing/uniform) and is trained by minimizing KL between categorical true posteriors $q(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x}_{0})$ and $q_{\psi}(\mathbf{z}_{s}\mid\mathbf{z}_{t})$.
+We introduce Guided Transfer Learning (GTL), a transfer framework for discrete diffusion models that adapts a pretrained source model to a target domain without updating the denoiser.
+Our approach is orthogonal to fine-tuning: rather than updating the denoiser, GTL adapts sampling by reweighting the source reverse transitions with a learned density-ratio guidance model.
+In this section, we first present the core transfer identity for a single denoising step, then extend it to sequences and develop an efficient guided sampler.
+Why not fine-tune?
+Fine-tuning adapts the source denoiser $p_{\theta}$ to the target domain by optimizing the diffusion objective (e.g., equation 3) on target samples.
+This optimization becomes expensive for large language models and tends to overfit when target data are scarce, often degrading diversity.
+These small-data failures motivate our approach: rather than updating millions of denoiser parameters, we reuse $p_{\theta}$ as-is and place the adaptation burden on a much smaller auxiliary model.
+Key Assumption.
+This auxiliary model is meaningful only if the two domains share the same corruption mechanism.
+Let $p_{X}$ denote the source data distribution and $q_{X}$ the target data distribution over $\mathbf{x}_{0}$.
+We assume a shared forward process, $p(\mathbf{z}_{t}\mid\mathbf{x}_{0})=q(\mathbf{z}_{t}\mid\mathbf{x}_{0})$ for all $t$, which holds whenever both domains use the same discrete forward kernel (e.g., an identical masking schedule or transition family).
+Under this shared corruption, differences between the domains can be expressed at the data level through the density ratio $r(\mathbf{x}_{0})\coloneqq q_{X}(\mathbf{x}_{0})/p_{X}(\mathbf{x}_{0}),$ assumed well-defined on the support of interest.
+This ratio quantifies how much more (or less) likely a clean sequence is under the target distribution than under the source, and it is exactly the quantity needed to reweight source reverse transitions into target reverse transitions.
+The next theorem makes this reweighting precise for a single reverse transition: it shows that the reverse kernel minimizing the target KL objective can be written as the source reverse kernel multiplied by a ratio-dependent weight.
 Theorem 1
@@
 $q_{\psi^{\star}}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})=\frac{p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}{\displaystyle\sum_{\mathbf{\tilde{z}}_{s(i)}}p(\mathbf{\tilde{z}}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{\tilde{z}}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}$ (6)
-In (7), $p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\approx p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ is the trained denoiser on the source domain, and the expectation factor $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\tfrac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right]$ is the guidance ratio.
-Thus, the target diffusion process can be approximated with two networks $\psi^{\star}=[\theta,\phi]$.
-The proof of this theorem can be found in the supplementary materials.
-The advantage of this approach is that it does not require fine-tuning the pretrained discrete diffusion model.
-Instead, the main computational cost is shifted to training a domain classifier using both domains as data.
-Importantly, the domain classifier can be much smaller than the discrete diffusion model.
-Moreover, density ratio estimation provides more flexibility in regularization when dealing with highly imbalanced data.
-It only requires distinguishing domains rather than modeling the full target distribution, which makes the training dynamics between the discriminative and generative models more stable.
-The result in (ouyang2024transferlearningdiffusionmodels) is derived only for score-based diffusion in continuous data domains.
-But our formulation in (7) offers two advantages: (i) it permits flexible definitions of the forward and reverse discrete diffusion processes, and (ii) the guidance term can be trained under arbitrary noise processes, in either continuous- or discrete-time settings.
-In the supplementary materials (Theorem 3), we extend the result to the score-based, continuous-time diffusion setting over discrete state spaces.
-The key mathematical step is to trace back through to the ELBO formulation used in (6).
-For the masked (absorbing-state) diffusion training objective, this link is directly shown in Equation (4).
-Extension to Sequence
-Leveraging our theoretical results, (7) closely parallels (5).
-Instead of a classifier $p\!\left(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)}\right)$, we use a ratio network $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\tfrac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right]$.
-To extend this to sequences, define $\mathbf{z}_{t,s^{\ell}}^{(1:L)}\coloneqq[\,\mathbf{z}_{t}^{(1:\ell-1)},\,\mathbf{z}_{s}^{(\ell)},\,\mathbf{z}_{t}^{(\ell+1:L)}\,]$, i.e., the sequence identical to $\mathbf{z}_{t}^{(1:L)}$ except at position $\ell$, where the token is replaced by $\mathbf{z}_{s}^{(\ell)}$.
-This yields the following factorized, ratio-guided conditional (with guidance strength $\gamma\geq 0$):
-$q_{\theta,\phi}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)}\right)$
-$=\prod_{\ell=1}^{L}\frac{\Big[r_{\phi}\!\big(\mathbf{z}_{t,s^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\displaystyle\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}\Big[r_{\phi}\!\big(\mathbf{z}_{t,\tilde{s}^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}.$ (8)
-A naive integration of this guidance into the sampling process incurs a computational complexity of $\mathcal{O}(|\mathcal{V}|\cdot L)$ per denoising step, since we must evaluate $r_{\phi}$ for all sequences obtained by replacing every possible token in $\mathcal{V}$ at each position $L$ in the denominator.
-Prior work (schiff2025simpleguidancemechanismsdiscrete; nisonoff2025unlockingguidancediscretestatespace) targets settings with modest vocabulary sizes and sequence lengths, where this cost is manageable.
-In our experiments with the Markov chain (Section 5.1) it is also not prohibitive.
-But for text applications, the cost becomes significant.
-(nisonoff2025unlockingguidancediscretestatespace) proposes a first-order Taylor approximation that reduces evaluations, but it is accurate primarily in the high-noise regime and degrades as $t$ decreases.
-4.1 Scalability via Planner Sampling
-In this section, we tackle the main bottleneck: integrating guidance efficiently at inference time. The gains from a set of simple techniques are summarized in Table 1.
+Practical implementation.
+Equation 6 (proved in Appendix C) expresses the target reverse transition as a reweighted version of the source reverse transition.
+This expression suggests a direct implementation with two components: a frozen source denoiser that provides $p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ and a learned ratio module that provides the ratio-dependent weight.
+Concretely, we approximate the source kernel with the pretrained denoiser $p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$, and we learn a ratio function $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right].$ Together, these two networks define the GTL sampler, $\psi^{\star}=[\theta,\phi]$, which generates target-domain samples while keeping the denoiser fixed.
+Why ratio guidance helps in the low-data regime?
+The key advantage of equation 6 is that it eliminates the need to fine-tune the denoiser on scarce target data.
+Instead, adaptation is shifted to learning a domain discriminator (or an equivalent ratio estimator) from mixed source/target samples, which can be orders of magnitude smaller than the diffusion model.
+This shift is especially helpful under strong source–target imbalance: ratio estimation only needs to separate domains, rather than model the full target distribution, and it admits flexible regularization that tends to stabilize training in data-scarce settings.
+Relation to TLDM and generality of our formulation.
+TLDM (Ouyang et al., 2024) derives a ratio-transfer rule only for continuous data score-based diffusion in $\mathbb{R}^{d}$, where guidance is expressed through Euclidean score gradients such as $\nabla_{x_{t}}\log q_{t}(x_{t})$.
+These gradients are not available for categorical tokens, and discrete diffusion is commonly trained by minimizing KL divergences between categorical posteriors rather than score-matching MSE.
+Our derivation therefore works directly with discrete forward and reverse kernels and the ELBO-style objective in equation 5.
+This formulation is architecture- and corruption-agnostic: it applies to a range of discrete forward processes (e.g., uniform, absorbing/masked, and general D3PM transitions) and does not depend on a particular parameterization of the denoiser in benefit to TLDM.
+Continuous-Time GTL for Discrete Diffusion.
+Because the argument is expressed at the level of the ELBO/KL objective, the same ratio-transfer principle carries over to continuous-time discrete diffusion whenever the training loss can be related back to an ELBO (e.g., equation 3 and equation 36).
+In Appendix (Theorem 3), we instantiate this transfer rule for continuous-time, score-based diffusion over discrete state spaces by reweighting the reverse-rate matrix with the learned ratio.
+Sequence Formulation.
+We now apply Theorem 1 to length-$L$ token sequences.
+As in prior work on discrete classifier guidance, we adopt a token-factorized reverse transition, which yields a practical sampler while keeping the denoiser unchanged.
+Concretely, we replace the classifier term $p(y\mid\mathbf{z}_{s}^{(\ell)},\mathbf{z}_{t}^{(1:L)})$ in equation 4 with a learned ratio model.
+In the sequence setting, the ratio model is evaluated on a single-token intervention of the current noisy sequence.
+Define $\mathbf{z}_{t,s^{\ell}}^{(1:L)}\coloneqq[\,\mathbf{z}_{t}^{(1:\ell-1)},\,\mathbf{z}_{s}^{(\ell)},\,\mathbf{z}_{t}^{(\ell+1:L)}\,]$, i.e., the sequence obtained by replacing position $\ell$ in $\mathbf{z}_{t}^{(1:L)}$ by a candidate token $\mathbf{z}_{s}^{(\ell)}$.
+With guidance strength $\gamma\geq 0$, the resulting ratio-guided, factorized reverse transition is equation 7.
+$q_{\theta,\phi}\!\left(\mathbf{z}_{s}^{(1:L)}\mid\mathbf{z}_{t}^{(1:L)}\right)=\prod_{\ell=1}^{L}\frac{\Big[r_{\phi}\!\big(\mathbf{z}_{t,s^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\mathbf{z}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}{\displaystyle\sum_{\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}}\Big[r_{\phi}\!\big(\mathbf{z}_{t,\tilde{s}^{\ell}}^{(1:L)}\big)\Big]^{\gamma}\;p_{\theta}\!\left(\tilde{\mathbf{z}}_{s}^{(\ell)}\mid\mathbf{z}_{t}^{(1:L)}\right)}.$ (7)
+A naive implementation of equation 7 is prohibitively expensive.
+For each masked position $\ell$, the denominator requires evaluating $r_{\phi}(\mathbf{z}_{t,\tilde{s}^{\ell}}^{(1:L)})$ for all $\tilde{\mathbf{z}}_{s}^{(\ell)}\in\mathcal{V}$, resulting in $\mathcal{O}(L|\mathcal{V}|)$ ratio evaluations per denoising step.
+This is manageable for small $L$ or $|\mathcal{V}|$ (e.g., our Markov-chain setting), but becomes prohibitive for text.
 Method Per-Step Complexity Number of Steps ($T$)
@@
 Table 1: Computational cost per denoising step and total number of steps.
-Ancestral Sampling
-In the naive implementation, at all denoising steps, we compute logits from the denoiser and from the guidance network with $\mathcal{O}(|\mathcal{V}|\cdot L)$ complexity.
-Following (sahoo2024simpleeffectivemaskeddiffusion), in masked diffusion, we can remove explicit time conditioning because the number of masked tokens in the sequence already implies the timestep.
-This is not possible for uniform noise, which mixes token identities.
-We therefore cache logits across steps.
-If a step does not change the sequence, we reuse the logits from the previous step.
-The number of sampling steps where the network is called then becomes upper-bounded by the sequence length, since in the worst case, every position unmasks in a different step.
-In addition, we introduce a top-$n$ style pruning mechanism for the ratio network.
-Specifically, we assume that (i) the denoiser distribution $p_{\theta}(\cdot\mid z_{t})$ is sufficiently concentrated, so that the true target distribution $q$ differs from $p$ only within the top-$n_{\text{ratio}}$ candidates, and (ii) the density ratio $r(z)=q(z)/p(z)$ is bounded, in the sense that there exists a constant $C<\infty$ such that for all tokens within the top-$n_{\text{ratio}}$ set of $p_{\theta}$, we have $\frac{q(z)}{p(z)}\leq C$.
-Outside this set, where $p(z)$ is already very small, no such bound is required.
-Consequently, we evaluate the ratio term only for the top $n_{\text{ratio}}$ candidates proposed by the denoiser at each position.
-This results in position-specific candidate sets in the denominator of equation (8) and reduces the computational cost of ratio evaluations to $\mathcal{O}(n_{\text{ratio}}\cdot L)$ with $n_{\text{ratio}}\leq|\mathcal{V}|$.
-In the experiments, the combination of top-$n_{\text{ratio}}$ selection and the guidance weight $\gamma$ can misallocate probability mass, which may cause the diffusion process to collapse into a fully masked sequence at the final step.
-Thus, a key assumption is that the probability of remaining a masked token must be equal between the target and source distributions.
-As a result, we enforce a position-independent normalization which implies the renormalized update over non-mask tokens, thus stabilizing the denoising diffusion process.
-Further details of this stabilizer are presented in the supplementary materials.
 Algorithm 1 Ratio-Guided Denoise Step $(\mathbf{z}_{t},t,\Delta t)$.
@@
 13:return $\mathbf{z}_{s}$
+3.1 Scalability via Planner Sampling
+Our goal is to reduce ratio-network evaluations from $\mathcal{O}(L|\mathcal{V}|)$ to a budget that is independent of $|\mathcal{V}|$ and ideally weakly dependent on $L$.
+Table 1 summarizes how caching, top-$n_{\text{ratio}}$, and planner sampling progressively reduce this cost.
+Ancestral Sampling
+In the naive implementation, at all denoising steps, we compute logits from the denoiser and from the guidance network with $\mathcal{O}(|\mathcal{V}|\cdot L)$ complexity.
+Caching.
+In masked diffusion, the timestep is implicitly determined by the number of masked tokens, allowing us to remove explicit time conditioning from the networks (Sahoo et al., 2024).
+As a result, if a denoising step does not modify the sequence, the corresponding logits can be reused.
+By caching logits across steps, the number of network evaluations becomes upper-bounded by the number of unmasking events.
+In the worst case, each position unmasks once, yielding at most $L$ effective network calls.
+This optimization does not apply to uniform-noise diffusion, where token identities are mixed.
+Top-$n_{\text{ratio}}$ pruning.
+Although caching reduces the number of steps, each step still requires evaluating the ratio network over the full vocabulary.
+To address this, we restrict ratio evaluations to a small candidate set.
+We assume that the denoiser distribution $p_{\theta}(\cdot\mid\mathbf{z}_{t})$ is sufficiently concentrated, so that meaningful deviations between the source and target distributions occur primarily among the top-$n_{\text{ratio}}$ tokens.
+Formally, we assume that for these candidates the density ratio $r(z)=q(z)/p(z)$ is bounded by a constant $C<\infty$; no such bound is required outside this set, where $p(z)$ is already negligible.
+Under this assumption, we evaluate the ratio term only for the top-$n_{\text{ratio}}$ tokens proposed by the denoiser at each position.
+This produces position-specific candidate sets in the denominator of equation 7 and reduces the cost of ratio evaluations to $\mathcal{O}(n_{\text{ratio}}\,L)$, with $n_{\text{ratio}}\ll|\mathcal{V}|$ in practice.
+Mask-probability stabilizer.
+The interaction between top-$n_{\text{ratio}}$ pruning and the guidance weight $\gamma$ can distort probability mass, occasionally causing the diffusion process to collapse to fully masked sequences.
+To prevent this failure mode, we enforce a position-independent constraint that preserves the probability of remaining masked across source and target distributions.
+This induces a renormalized update over non-mask tokens and stabilizes the denoising dynamics.
+Details of this stabilization are provided in the appendix.
 Planner Sampling
-We also apply our method to text in Section 5.2, where $n_{\text{ratio}}\ll\lvert\mathcal{V}\rvert$, yet sampling remains slow and costly mainly because of the $L$ in the complexity.
-To address this, we propose a planner-based algorithm.
-The planner $\rho_{\vartheta}$ is trained to predict the next denoising position in the masked sequence.
-Concretely, we replace the Gillespie process with $\tau$-leaping by a planner schedule that unmasks exactly one position per step, thereby fixing the number of denoising steps to $L$.
-In many implementations, $T\approx L$, so enforcing a hard budget of $L$ steps does not materially change the step count, while reducing guidance evaluations and wall-clock time.
-Pseudocode is provided in Algorithm 1.
-An important achievement here is that the loop in lines 4 to 7 can now be parallelized by vectorization on a single GPU.
-As a result, only one forward pass of the three networks is required.
-For training the planner, we adopt a simplified training algorithm following (peng2025pathplanningmaskeddiffusion; liu2025thinkgeneratediscretediffusion).
-For ratio training, we use a discrete data training loss function building on the work of (ouyang2024transferlearningdiffusionmodels).
-Further details of the loss function and training algorithm are provided in the supplementary materials.
+Even with top-$n_{\text{ratio}}$ pruning, the overall complexity still scales linearly with $L$.
+To further reduce cost, we introduce a planner-based sampler that selects a single position to update at each step.
+The planner $\rho_{\vartheta}$ is trained to predict the next denoising position in the masked sequence, replacing the stochastic Gillespie process with a deterministic $\tau$-leaping schedule.
+This strategy fixes the number of denoising steps to exactly $L$, which is comparable to standard masked diffusion where typically $T\approx L$, while substantially reducing guidance evaluations and wall-clock time.
+Algorithm 1 summarizes the resulting ratio-guided denoising step.
+Crucially, once the planner selects a position, ratio evaluations over the top-$n_{\text{ratio}}$ candidates can be fully vectorized.
+As a result, each denoising step requires only a single forward pass through the denoiser, ratio network, and planner.
+For planner training, we follow simplified strategies from Peng et al.
+(2025); Liu et al.
+(2025).
+The ratio network is trained using a discrete-data loss adapted from Ouyang et al.
+(2024).
+Additional training details are provided in the supplementary materials.
 4 Experiments
@@
 Sampling from the source domain proceeds in the same way.
-Implementation
-We use the sampling process of (sahoo2024simpleeffectivemaskeddiffusion) for the source, finetuned, and target discrete diffusion models.
-The backbone is a Diffusion Transformer (DiT) (peebles2023scalablediffusionmodelstransformers) with about five million learnable parameters.
-The ratio estimator and the classifier use about half as many parameters with a final mean and fully connected layer that maps the output to a scalar.
-The exact implementation appears in our codebase.
-For guided sampling, we follow the ancestral sampling method.
-Since the vocabulary is small we set $n_{\text{ratio}}=N$ and use $20$ sampling steps.
-We train a masked diffusion model on the source domain for $30$ epochs.
-We then compare a vanilla masked diffusion model trained directly on the target samples (60 epochs), a finetuned diffusion model (90 epochs), and our GTL method (60 epochs).
-The finetuned model and our method share the same source diffusion model.
+Implementation details can be found in the appendix E.1.
 Results
@@
 In this section, we apply our method to text data.
-We use the arXiv abstracts dataset (arxiv_org_submitters_2024), which contains 1.7 million arXiv articles with titles, authors, categories, and abstracts.
+We use the arXiv abstracts dataset (arXiv.org submitters, 2024), which contains 1.7 million arXiv articles with titles, authors, categories, and abstracts.
 We focus on Computer Science, Mathematics, and Physics.
@@
 Implementation and Evaluations
-We train the source diffusion model for 100,000 gradient steps using the DiT architecture (peebles2023scalablediffusionmodelstransformers) with 59.8M trainable parameters.
+We train the source diffusion model for 100,000 gradient steps using the DiT architecture (Peebles and Xie, 2023) with 59.8M trainable parameters.
 A log–linear noise scheduler is used for both training and sampling.
@@
 For each method we generate 128 sequences of length 512.
-We evaluate quality with MAUVE (pillutla2021mauvemeasuringgapneural), which captures the tradeoff between perplexity and diversity.
+We evaluate quality with MAUVE (Pillutla et al., 2021), which captures the tradeoff between perplexity and diversity.
 Although generative perplexity (Gen.
-PPL) is often reported, we observed, consistent with prior work (wang2025remaskingdiscretediffusionmodels), that Gen.
+PPL) is often reported, we observed, consistent with prior work (Wang et al., 2025), that Gen.
 PPL can remain low even when the text has reduced utility, such as repeated words.
@@
 The classifier follows the setup used for the ratio model in the supplementary materials with label smoothing set to $0.05$.
-Behavior of $\gamma$ and $n_{\text{ratio}}$
+4.2.1 Behavior of $\gamma$ and $n_{\text{ratio}}$
 We study the sensitivity of GTL to the guidance weight $\gamma$ and the candidate budget $n_{\text{ratio}}$.
@@
 Overall, the results suggest that $n_{\text{ratio}}\approx 256\!-\!512$ and $\gamma\approx 4\!-\!6$ achieve a favorable balance, providing strong target alignment with minimal quality loss.
-Method Comparison and Sensitivity to Data Fraction
+4.2.2 Method Comparison and Sensitivity to Data Fraction
 We compare three methods: a fine-tuned discrete diffusion model trained on the target domain after pretraining on the source domain, a vanilla discrete diffusion model trained directly on the target domain, and our Guided Transfer Learning (GTL) method with the ratio estimator trained across target-data fractions of 100%, 30%, 5%, and 1%.
@@
 We demonstrate the effectiveness of our method on synthetic Markov chains and arXiv abstracts (language modeling), particularly when the target dataset is small relative to the source.
-Limitations
-Overall, the proposed method is promising and applicable to a wide range of tasks involving domain shifts.
+Limitations Overall, the proposed method is promising and applicable to a wide range of tasks involving domain shifts.
 However, it requires access to both source and target domains.
@@
 It would also be interesting to evaluate the learned ratio on quality-filtered tasks, such as reducing toxicity in language modeling.
-7 Supplementary Materials for Guided Transfer Learning for Discrete Diffusion Models
-Appendix A Proof of Theorem 1
+Appendix A Appendix
+Appendix B Related Work
+Discrete Diffusion Models
+Discrete diffusion models (DDMs) have become a standard choice for generative modeling over categorical sequences, including biological sequences (DNA, RNA, proteins) and text (Sarkar et al., 2024; Yu et al., 2025).
+Compared to approaches that add continuous noise to embeddings or latents, fully discrete formulations avoid the projection step back to tokens, which can introduce an embedding-to-token mismatch (Li et al., 2022b; Lovelace et al., 2023).
+Two families of discrete diffusion methods are especially common.
+The first is discrete-time diffusion in the D3PM framework, where a Markov corruption process is defined by transition matrices and a denoiser is trained to invert the forward kernels; simplified masked variants such as MLDM and MD4 fall into this category (Sahoo et al., 2024; Shi et al., 2025; Austin et al., 2023).
+The second is continuous-time discrete diffusion, which parameterizes a discrete score (a log-ratio between neighboring-time marginals) and induces a continuous-time Markov chain (CTMC) for generation (Lou et al., 2024; Meng et al., 2023; Campbell et al., 2022).
+Guided Diffusion Models
+Guidance is a widely used mechanism for steering diffusion sampling in both continuous and discrete domains.
+In discrete diffusion, classifier-free guidance typically combines conditional and unconditional predictions at inference time, requiring two forward passes of a single model trained on a mixture of conditional and unconditional examples (Huang and Tang, 2025).
+Classifier-based guidance is closer to our setting: it introduces an auxiliary classifier $p(y\mid x_{t},t)$ and reweights the reverse transition so that samples are biased toward the desired attribute y, while the denoiser itself is trained on the full corpus (Schiff et al., 2025; Nisonoff et al., 2025).
+A direct implementation scores token substitutions across all positions and vocabulary items, requiring $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies.
+Because each evaluation corresponds to a distinct single-token intervention on the full sequence, this cost is not removed by straightforward parallelization.
+To reduce this cost, (Nisonoff et al., 2025) proposes a first-order Taylor approximation that replaces many evaluations with a forward and backward pass.
+Related alternatives adjust sampling using inference-time optimization: TESS2 (Tae et al., 2025) performs gradient ascent on a reward with respect to the model logits, and energy-based diffusion language models define an energy over predicted clean sequences and use it to reweight denoising updates to reduce train–sample mismatch (Xu et al., 2025).
+Transfer Learning
+Fine-tuning and transfer learning share the goal of adapting a model to a new task or domain.
+Fine-tuning is computationally expensive because it updates the pretrained denoiser’s weights on new data.
+Prior work on continuous diffusion models explores reinforcement learning (Fan et al., 2023; Clark et al., 2024) and adapter-based fine-tuning (Moon et al., 2022; Xie et al., 2023a).
+In contrast, transfer learning for diffusion remains comparatively underexplored.
+Transfer Learning for Diffusion Models (TLDM) (Ouyang et al.
+(2024)) proposes a ratio-estimation technique for continuous score-based models,but it relies on continuous scores and is evaluated in low-dimensional settings.
+Our work closes the gap for discrete diffusion by deriving ratio-guided reverse transitions in both discrete-time, continuous-time and score based formulations for high dimansion application.
+Appendix C Proof of Theorem 1
 For the proof, we use the formulation of importance reweighting as described in the following lemma.
@@
 $=\sum_{i=1}^{T}\mathbb{E}_{\mathbf{x}_{0}\sim p}\;\mathbb{E}_{\mathbf{z}_{t(i)}\sim p(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}\Bigl[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}D_{\mathrm{KL}}\!\bigl(p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})\,\|\,q_{\psi}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\bigr)\,\Bigr]$ (14)
-To conclude Lemma 2, we leverage the fact that the source and target domain share the same forward kernel in Equation 14 to 15 $q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=\frac{q(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\cdot q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})}{q(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$.
+To conclude Lemma 2, we leverage the fact that the source and target domain share the same forward kernel in Equation 13 to 14 $q(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=\frac{q(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\cdot q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})}{q(\mathbf{z}_{t(i)}\mid\mathbf{x}_{0})}=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$.
 Now, we use Lemma 2:
@@
 $={\frac{p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}{\displaystyle\sum_{\mathbf{\tilde{z}}_{s(i)}}p(\mathbf{\tilde{z}}_{s(i)}\mid\mathbf{z}_{t(i)})\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{\tilde{z}}_{s(i)})}\Bigr[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\Bigl]}}$ (28)
-For 27 to 28 we show that $\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}$:
+For 26 to 27 we show that $\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})=p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}$:
 $\displaystyle\sum_{\mathbf{x}_{0}}\alpha(\mathbf{x}_{0},\mathbf{z}_{t(i)})\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$ $=\displaystyle\sum_{\mathbf{x}_{0}}p(\mathbf{x}_{0}\mid\mathbf{z}_{t(i)})\,\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\,p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)},\mathbf{x}_{0})$ (29)
@@
 $=\frac{p(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})}{p(\mathbf{z}_{t(i)})}\displaystyle\sum_{\mathbf{x}_{0}}q(\mathbf{x}_{0})p(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})=p(\mathbf{z}_{t(i)}\mid\mathbf{z}_{s(i)})\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{t(i)})}$ (31)
-For 28 to 29 we show that $\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}=\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\bigl[q(\mathbf{x}_{0})/p(\mathbf{x}_{0})\bigr]$:
+For 27 to 28 we show that $\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}=\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\bigl[q(\mathbf{x}_{0})/p(\mathbf{x}_{0})\bigr]$:
 $\frac{q(\mathbf{z}_{s(i)})}{p(\mathbf{z}_{s(i)})}$ $=\frac{\sum_{\mathbf{x}_{0}}q(\mathbf{z}_{s(i)}\mid\mathbf{x}_{0})\,q(\mathbf{x}_{0})}{p(\mathbf{z}_{s(i)})}$ (32)
@@
 $=\sum_{\mathbf{x}_{0}}p(\mathbf{x}_{0}\mid\mathbf{z}_{s(i)})\;r({\mathbf{x}_{0}})=\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\bigl[q(\mathbf{x}_{0})/p(\mathbf{x}_{0})\bigr]$ (34)
-Appendix B Guided $\tau$-Sampling
+Appendix D Guided $\tau$-Sampling
 Algorithm 2 Naive Guided Denoise Step $(\mathbf{z}_{t},\ t,\ \Delta t)$
@@
 We use this algorithm in the first experiments section.
-Appendix C Training Details
+Appendix E Training Details
 The Code will be released upon acceptance.
@@
 11:return trained planner parameters $\vartheta$
-Appendix D Application to continuous-time Score-based Models
-D.1 Background
-(campbell2022continuoustimeframeworkdiscrete) introduces a continuous-time framework for discrete denoising models formulated as a continuous-time Markov chain (CTMC).
-Unlike the discrete-time setting of (austin2023structureddenoisingdiffusionmodels) with steps $t=0,1,\dots,T$, here $t\in[0,T]$ is continuous.
+E.1 Implementation on Markov Chain
+We use the sampling process of (Sahoo et al., 2024) for the source, finetuned, and target discrete diffusion models.
+The backbone is a Diffusion Transformer (DiT) (Peebles and Xie, 2023) with about five million learnable parameters.
+The ratio estimator and the classifier use about half as many parameters with a final mean and fully connected layer that maps the output to a scalar.
+The exact implementation appears in our codebase.
+For guided sampling, we follow the ancestral sampling method.
+Since the vocabulary is small we set $n_{\text{ratio}}=N$ and use $20$ sampling steps.
+We train a masked diffusion model on the source domain for $30$ epochs.
+We then compare a vanilla masked diffusion model trained directly on the target samples (60 epochs), a finetuned diffusion model (90 epochs), and our GTL method (60 epochs).
+The finetuned model and our method share the same source diffusion model.
+Appendix F Application to continuous-time Score-based Models
+F.1 Background
+Campbell et al.
+(2022) introduces a continuous-time framework for discrete denoising models formulated as a continuous-time Markov chain (CTMC).
+Unlike the discrete-time setting of (Austin et al., 2023) with steps $t=0,1,\dots,T$, here $t\in[0,T]$ is continuous.
 In this section, we switch notation.
@@
 We can express the reverse process in terms of the forward transition rates and a concrete score defined by $s_{\theta}(\mathbf{y})_{\mathbf{y}^{\prime}}\approx\tfrac{q_{t}(\mathbf{y}^{\prime}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}$, parameterized by a neural network $s_{\theta}:\mathcal{V}\to\mathbb{R}_{>0}^{|\mathcal{V}|}$.
-This enables training via concrete score matching (CSM) (meng2023concretescorematchinggeneralized), directly analogous to score-based models in continuous domains (song2021scorebasedgenerativemodelingstochastic).
-(lou2024discretediffusionmodelingestimating) propose the diffusion-weighted denoising score entropy (DWDSE) loss (their CDM-based objective) and note it can also be derived from the continuous-time likelihood framework of (campbell2022continuoustimeframeworkdiscrete); a detailed derivation appears in Appendix C.3 of (sahoo2024simpleeffectivemaskeddiffusion).
+This enables training via concrete score matching (CSM) (Meng et al., 2023), directly analogous to score-based models in continuous domains (Song et al., 2021).
+Lou et al.
+(2024) propose the diffusion-weighted denoising score entropy (DWDSE) loss (their CDM-based objective) and note it can also be derived from the continuous-time likelihood framework of (Campbell et al., 2022); a detailed derivation appears in Appendix C.3 of (Sahoo et al., 2024).
 Concretely,
@@
 where $K(a)=a\log a-a$.
-D.2 Theorem
+F.2 Theorem
 In the previous section, we introduced the diffusion process and the source-domain objective on which the score function $s_{\theta}$ is trained.
-We now state a theorem that enables sampling from the target distribution without retraining $s_{\theta}$ under the objective in 37.
+We now state a theorem that enables sampling from the target distribution without retraining $s_{\theta}$ under the objective in 36.
 Let $r_{\phi}(\mathbf{y})$ denote the learned density-ratio guidance term defined earlier.
@@
 $\widetilde{R}_{t}^{\psi}(\mathbf{y}^{\prime},\mathbf{y})=\begin{cases}\dfrac{r_{\phi}(\mathbf{y}^{\prime})}{r_{\phi}(\mathbf{y})}\,\widetilde{R}_{t}^{\theta}(\mathbf{y}^{\prime},\mathbf{y}),&\mathbf{y}^{\prime}\neq\mathbf{y},\\[8.0pt] -\displaystyle\sum_{\tilde{\mathbf{y}}\neq\mathbf{y}}\dfrac{r_{\phi}(\tilde{\mathbf{y}})}{r_{\phi}(\mathbf{y})}\,\widetilde{R}_{t}^{\theta}(\tilde{\mathbf{y}},\mathbf{y}),&\mathbf{y}^{\prime}=\mathbf{y}.\end{cases}$ (37)
-D.3 Proof of Theorem 3
-For the proof, we take the loss for a small step $\Delta t$ with $s=t-\Delta t$ from (36) and apply Theorem 1, which yields the following expression:
+F.3 Proof of Theorem 3
+For the proof, we take the loss for a small step $\Delta t$ with $s=t-\Delta t$ from equation 35 and apply Theorem 1, which yields the following expression:
 $q_{s\mid t}^{\psi^{\star}}(\mathbf{y}^{\prime}\mid\mathbf{y})$ $=\frac{p_{s\mid t}(\mathbf{y}^{\prime}\mid\mathbf{y})\,\mathbb{E}_{\mathbf{x}\sim p(\cdot\mid\mathbf{y}^{\prime})}\!\bigl[\tfrac{q(\mathbf{x})}{p(\mathbf{x})}\bigr]}{\displaystyle\sum_{\widetilde{\mathbf{y}}}p_{s\mid t}(\widetilde{\mathbf{y}}\mid\mathbf{y})\,\mathbb{E}_{\mathbf{x}\sim p(\cdot\mid\widetilde{\mathbf{y}})}\!\bigl[\tfrac{q(\mathbf{x})}{p(\mathbf{x})}\bigr]}$ (38)
@@
 […]
-The work from (nisonoff2025unlockingguidancediscretestatespace) in equation (2) arrives at a similar term just for classifier-based guidance.
+The work from (Nisonoff et al., 2025) in equation (2) arrives at a similar term just for classifier-based guidance.

### v4

@@
 (2025) have explored CBG mechanisms for discrete DMs.
-These methods apply mainly to domains with small $|\mathcal{V}|$ and modest $L$, such as biological sequences, leaving scalable CBG for language modeling as an open problem.
+These methods apply mainly to domains with small vocabulary size $|\mathcal{V}|$ and modest sequence length $L$, such as biological sequences, leaving scalable CBG for language modeling as an open problem.
 This constaints are due to the necessity of token substitutions across all positions and vocabulary items during score computation.
-More precisely, letting $L$ be the sequence length and $\mathcal{V}$ the vocabulary, naive CBG requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies.
+More precisely, naive CBG requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies.
 In this paper, we aim to leverage CBGs for transfer learning, while overcoming the limiting small vocabulary sizes.
@@
 The forward process generates $\mathbf{z}_{t}$ by injecting noise into clean data.
-We consider discrete latents: $\mathbf{z}_{t}\in\mathcal{V}$, the set of one-hot tokens over a vocabulary of size $N$, $\mathcal{V}\coloneqq\{\mathbf{z}\in\{0,1\}^{N}:\sum_{i=1}^{N}z_{i}=1\}\subset\Delta^{N-1},$ where $\Delta^{N-1}$ denotes the probability simplex over $N$ categories.
+We consider discrete latents: $\mathbf{z}_{t}\in\mathcal{V}$, the set of one-hot tokens over a vocabulary $\mathcal{V}$ of size $N$, $\mathcal{V}\coloneqq\{\mathbf{z}\in\{0,1\}^{N}:\sum_{i=1}^{N}z_{i}=1\}\subset\Delta^{N-1},$ where $\Delta^{N-1}$ denotes the probability simplex over $N$ categories.
 Let $\text{Cat}(\cdot;\mathbf{\pi})$ be the categorical distribution over $N$ classes with the probability vector $\mathbf{\pi}\in\Delta^{N-1}$.
@@
 Concretely, we approximate the source kernel with the pretrained denoiser $p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$, and we learn a ratio function $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right].$ Together, these two networks define the GTL sampler, $\psi^{\star}=[\theta,\phi]$, which generates target-domain samples while keeping the denoiser fixed.
+The ratio network is trained using a discrete-data loss adapted from Ouyang et al.
+(2024).
+Additional training details are provided in the supplementary materials.
 Why ratio guidance helps in the low-data regime?
@@
 (2025).
-The ratio network is trained using a discrete-data loss adapted from Ouyang et al.
-(2024).
-Additional training details are provided in the supplementary materials.
 4 Experiments
@@
 Refer to caption (a) SourceRefer to caption (b) Target
-n Vanilla Discrete Diffusion Finetuned Discrete Diffusion GTL
+$\bm{n}$ Vanilla Discrete Diffusion Finetuned Discrete Diffusion GTL
 1000 Refer to caption Refer to caption Refer to caption
@@
 20 Refer to caption Refer to caption Refer to caption
-Figure 2: True source and target (left). Estimated target transition matrices for varying $n$ across three methods (right).
-Experimental Setup
-We construct two Markov chains with distinct transition matrices.
+Figure 2: True source and target (left). Estimated target transition matrices for varying number of target samples $n$ across three methods (right).
+Experimental Setup We construct two Markov chains with distinct transition matrices.
 Let the probability state vector be $\mathbf{\pi}_{i}\in\Delta^{N-1}$ over a vocabulary of size $N=5$.
@@
 Implementation details can be found in the appendix E.1.
-Results
-We consider $m=10000$ source samples and $n\in\{1000,100,20\}$ target samples.
+Results We consider $m=10000$ source samples and $n\in\{1000,100,20\}$ target samples.
 We estimate the transition matrix by sampling $4096$ sequences, counting transitions between states, and normalizing by the total count.
@@
 We observe that the vanilla target diffusion model and the finetuned model both degrade in performance when fewer target samples are available.
-$n_{\text{tgt}}$ Vanilla (Target-only) Finetuned GTL
+$n$ Vanilla (Target-only) Finetuned GTL
 […]
@@
 20 $0.5842\pm 0.0244$ $0.4004\pm 0.0983$ $\mathbf{0.3621}\pm 0.0478$
-Table 2: Comparison of baseline methods and best guided results (mean $\pm$ standard error; n=3n{=}3 seeds) across target sample sizes for $(\text{diag}_{\text{src}}=0.1,\text{diag}_{\text{tgt}}=0.8)$.
+Table 2: Comparison of baseline methods and best guided results (mean $\pm$ standard error; 33 seeds) across target sample sizes for $(\text{diag}_{\text{src}}=0.1,\text{diag}_{\text{tgt}}=0.8)$.
 Reported values are the average Kullback–Leibler divergence between true and estimated transition matrices; lower is better.
@@
 4.2 Language Modeling
-Refer to caption Refer to caption Refer to caption
-(a) Effect of $n_{\text{ratio}}$ at $\gamma=5$. (b) Effect of low $\gamma$ at fixed $n_{\text{ratio}}$ (c) MAUVE vs. high $\gamma$
 […]
@@
 We ensure that the three category subsets are disjoint.
-We define the union of Computer Science and Mathematics as the source domain and Physics as the target domain.
 We run two types of experiments.
@@
 The third is our guided transfer method.
-Implementation and Evaluations
-We train the source diffusion model for 100,000 gradient steps using the DiT architecture (Peebles and Xie, 2023) with 59.8M trainable parameters.
+Implementation and Evaluations We train the source diffusion model for 100,000 gradient steps using the DiT architecture (Peebles and Xie, 2023) with 59.8M trainable parameters.
 A log–linear noise scheduler is used for both training and sampling.
@@
 4.2.2 Method Comparison and Sensitivity to Data Fraction
+In this section, we define the union of Computer Science and Mathematics abstracts as the source domain and Physics abstracts as the target domain.
 We compare three methods: a fine-tuned discrete diffusion model trained on the target domain after pretraining on the source domain, a vanilla discrete diffusion model trained directly on the target domain, and our Guided Transfer Learning (GTL) method with the ratio estimator trained across target-data fractions of 100%, 30%, 5%, and 1%.
@@
 In contrast, the finetuned and vanilla discrete diffusion models show a stepwise drop in performance as the fraction decreases.
+Even in the full-target-data setting, GTL outperforms the baselines because there is still a large imbalance between the number of source-domain and target-domain samples.
 The GTL remains relatively stable, except in the 1% case, where we also observe a sharp decline.